### v1

# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

Abstract
Large-scale pre-training using videos has proven effective for robot learning. However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots. To remedy this, we propose H2R, a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric videos. This process explicitly bridges the visual gap between human and robot embodiments during pre-training. We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data. Based on this, we construct and release a family of 1M-scale datasets covering multiple robot embodiments (UR5 with gripper/Leaphand, Franka) and data sources (SSv2, Ego4D). To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions. We validate H2R across three simulation benchmarks—Robomimic, RLBench, and PushT and real-world manipulation tasks with a UR5 robot equipped with Gripper and Leaphand end-effectors. H2R consistently improves downstream success rates, yielding gains of 5.0%–10.2% in simulation and 6.7%–23.3% in real-world tasks across various visual encoders and policy learning methods. These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains.
Figure 1: Overview of H2R. H2R is a data augmentation technique designed to enhance robot pre-training by converting first-person human hand operation videos into robot-centric visual data. By bridging the visual domain gap, H2R improves pre-trained visual encoders for downstream robot policies (imitation learning), validated across simulation benchmarks and real-world robotic tasks.
1 Introduction
Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics. Inspired by the remarkable accomplishments of large scale pre-training in computer vision he2022mae ; nair2022r3m ; zeng2024learningmanipulationpredictinginteraction ; karamcheti2023voltron ; datacomp and natural language processing devlin2018bert ; yu2022pointbert ; openai2024gpt4technicalreport ; anil2023palm2technicalreport ; li2024datacomplm , many efforts have been devoted to harness large-scale data to construct generalizable representations in the robotics field embodimentcollaboration2024openxembodimentroboticlearning ; brohan2023rt1roboticstransformerrealworld ; brohan2023rt2visionlanguageactionmodelstransfer . Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive duan2023ar2 ; jang2022bczzeroshottaskgeneralization ; brohan2023rt1roboticstransformerrealworld ; khazatsky2024droidlargescaleinthewildrobot ; fang2023rh20tcomprehensiveroboticdataset ; gao2024efficientdatacollectionrobotic ; lee2021pickandplacetacklingroboticstacking ; herzog2023deeprlscalesorting ; kalashnikov2021mtoptcontinuousmultitaskrobotic ; duan2024manipulate ; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
Recent works nair2022r3m ; xiao2022mvp ; zeng2024learning analyze such egocentric human video datasets such as Ego4D grauman2022ego4d , SSv2 goyal2017something , and Epic Kitchens damen2018scaling with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations. However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems. Specifically, when utilizing the robot expert data to fine-tune the pre-trained robotic representations for downstream robotic tasks, the model has to learn to bridge the visual gap between the first-person human hand and the robots in addition to acquiring nuanced task-specific skills demonstrated in the robot expert data. This would result in increased complexity during the fine-tuning process and suboptimal performance.
To mitigate this issue, we propose H2R (as shown in Figure 1), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation. H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video. Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR pavlakos2023reconstructinghands3dtransformers to accurately detect the position and posture of the human hand in egocentric videos. Then, we simulate the same robot state in simulators to obtain the mask of robot actions. In the second stage, we use the Segment Anything Model kirillov2023segment to automatically separate human hand from background, and use the inpainting model LaMa suvorov2021resolutionrobustlargemaskinpainting to fill the removed hand mask. After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video. Finally, we overlay the robotic arm images captured by the simulator’s camera onto the areas where the human hands are removed. Through such a process, H2R explicitly reduces the gap between human and robot hands by creating realistic robotic arm movements that visually mimic human hand actions. It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
Based on this pipeline, we construct and release a series of large-scale robot-centric datasets (H2R-1M), each containing approximately 1 million augmented frames. These datasets cover multiple robot embodiments (UR5 with gripper/Leaphand, Franka) and human video sources (SSv2, Ego4D), and provide modular annotations to support downstream robotic learning. To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics. This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
To further verify the utility of the augmented datasets, we conduct downstream experiments comparing models pre-trained on original egocentric datasets with those trained on our released H2R-enhanced datasets. These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice. For the pre-training stage, we apply MAE he2022mae and R3M nair2022r3m frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video. The pretrained visual representations are used in downstream imitation learning pipelines. We conduct comprehensive evaluations across both simulation and real-world manipulation tasks. In simulation, we evaluate pre-trained visual encoders on three benchmark suites: Robomimic, RLBench, and PushT. H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, and +5.3% for MAE and +7.0% for R3M on the PushT task. In real-world experiments, we deploy H2R-enhanced encoders on a UR5 robot equipped with both Gripper and Leaphand end-effectors, evaluated over six manipulation tasks under two policy frameworks: Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act . H2R brings consistent improvements, with average success rate gains of +13.3% (MAE) and +17.0% (R3M) under DP, and +10.8% (MAE) and +17.5% (R3M) under ACT. H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively. These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulated and real-world robotic manipulation settings.
Our paper provides three contributions:
  • We propose a data-centric pipeline, H2R, to mitigate the gap between human and robot hands when utilizing large-scale egocentric video datasets to pre-train generalizable visual features for robots.
  • We construct and release diverse robot-centric datasets (H2R-1M), combining multiple embodiments and egocentric video sources to support robot-compatible visual pretraining.
  • We demonstrate the effectiveness of H2R through extensive experiments on closed-loop benchmarks.
2 Related Work
Robot Imitation Learning. Data-driven policy learning sergey2018learning ; luo2024hilserl ; xiao2022mvp ; yang2024equibotsim3equivariantdiffusionpolicy ; chi2023diffusionpolicy ; embodimentcollaboration2024openxembodimentroboticlearning has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation. Imitation learning chi2023diffusionpolicy ; ze2024dp3 ; zhao2023act ; mandlekar2021matters trains policies from successful demonstrations, often supervised by behavior cloning torabi2018behavioral ; florence2021implicit objectives. ACT zhao2023act addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models chi2023diffusionpolicy ; ze2024dp3 are introduced to handle the inherent multimodality of robot motions. In addition, CordViP fu2025cordvip leverages 3D object-robot correspondences to enhance dexterous manipulation.
Visual Encoder Pretraining for Robotics. Visual pretraining improves generalization of robotic policies across diverse tasks. Researchers have explored architectural designs He_2016_CVPR ; dosovitskiy12021an , training objectives He_2022_CVPR ; chen2020simple ; chen2020big , and dataset compositions deng2009imagenet ; lin2014microsoft ; shao2019object365 ; Schuhmann2021LAION400MOD . PVR-Control parisi22a shows that pretrained visual representations can outperform direct state-based policies. RPT radosavovic2023robot tokenizes observations to enable masked prediction pretraining. Methods like MVP radosavovic2022realworld and R3M nair2022r3m utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning. Voltron karamcheti2023voltron demonstrates the use of MAE and contrastive learning for hierarchical robot control.
Cross-Domain Visual Alignment. Bridging the domain gap between human and robot visual inputs remains a major challenge. WHIRL bahl2022humantorobotimitationwild matches task structure from third-person views, while RoVi-Aug chen2024roviaugrobotviewpointaugmentation and Mirage chen2024mirage manipulate appearance via segmentation or image-space preprocessing. EgoMimic kareer2024egomimicscalingimitationlearning removes hands and normalizes views to align egocentric perspectives.
3 H2R: Human-to-Robot Data Augmentation
In this section, we describe H2R, a data augmentation pipeline for robot learning from egocentric human videos (Figure 2). It replaces human hands in every frame with robotic arms equipped with various end effectors, generating a new, visually different dataset. This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains. Figure 3 illustrates examples of human hand videos that have been processed using H2R. These examples include different types of robots (UR5 and Franka) equipped with a variety of end-effectors, such as dexterous hands and grippers. We further introduce a CLIP-based metric to evaluate semantic consistency, and construct a family of large-scale datasets that combine different robot embodiments (UR5 with gripper, Leaphand, or both; Franka with gripper) and egocentric sources (SSv2 and Ego4D).
3.1 H2R Data Augmentation Pipeline
Figure 2: H2R Pipeline. H2R involves replacing human hands with robotic arms by first using the HaMeR model to detect hand poses and camera parameters. The human hand is then removed using the SAM, and the inpainting model LaMa fills in the gap. A robot hand is constructed based on the detected pose and keypoints, with the camera perspective adjusted to match the original image. Finally, the robot hand is overlaid onto the image, ensuring accurate alignment with the human hand.
3D Hand Pose Estimation. In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information. We adopt HaMeR pavlakos2023reconstructinghands3dtransformers , a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal. Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
Human Arm and Hand Removal. After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background. We use Segment Anything Model (SAM) kirillov2023segment to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR. Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa suvorov2021resolutionrobustlargemaskinpainting , a powerful inpainting model, to fill in the removed hand-arm region. This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
robotic arm and End Effector Construction. This step involves constructing the robotic arm and different end effectors. There are two common types of end-effectors in robotic manipulation: grippers and dexterous hands. In our pipeline, we handle them differently based on their structural characteristics. For dexterous hands, which have multiple degrees of freedom and resemble human hands, we compute the joint angles by analyzing the hand keypoints predicted by HaMeR. Specifically, each finger joint angle is determined by the angle formed between three consecutive keypoints corresponding to that finger segment. For grippers, which typically consist of two or more parallel fingers, we determine the degree of opening or closing based on the Euclidean distance between the relevant fingertips in the hand keypoints. However, since hand keypoints alone are insufficient to fully define the robotic arm’s complete configuration, especially for joints that do not directly correspond to hand movement, we manually assign plausible values to these undetermined joint positions to construct a reasonable arm pose.
Simulator Camera Position Alignment. The visual bias introduced by the camera perspective is more significant than the action retargeting itself. To address this, we use the hand keypoints and camera parameters from HaMeR to adjust the camera pose in the simulator. Specifically, we define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robotic arm in the simulator. By mapping the position of the camera in $C_{H}$ to $C_{S}$, we can ensure that the camera in the simulator shares the same perspective as the one captured in the real-world egocentric human image. The original camera position ${}^{W}\mathbf{cam}_{real}$ in the world frame is transformed to the aligned simulator position ${}^{W}\mathbf{cam}_{sim}$ using transformations from human hand (${}^{W}_{H}\mathbf{R}$) and robot simulator (${}^{W}_{S}\mathbf{R}$) coordinate systems:
${}^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (1)
A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix  B.
Robot Hand Rendering and Copy-paste. After setting the camera, the segmentation mask of the robotic arm is obtained by shooting with the camera. We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image. In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices. By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
3.2 Data Quality Evaluation
To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP radford2021learningtransferablevisualmodels . This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks. For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective. The human-centric template is defined as: A human is [action]”; the robot-centric template is defined as: A robotic arm is [action]” . Here, [action] is a natural language phrase describing the high-level behavior, such as holding a bottle” or “brushing paint on a wall.”
We use the CLIP ViT-B/32 model to compute cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt. This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description. Similarly, we compute the similarity between the original human frame and the human-centric prompt. This setup allows us to directly compare semantic alignment before and after augmentation under their respective modalities. For each frame, only the action-relevant region is retained, and background context is preserved via inpainting using LaMa. We average similarity scores over multiple samples per action class to mitigate variance and isolate the impact of robotic replacement. All textual prompts are manually curated to maintain consistency across frames and embodiments.
This evaluation protocol provides a lightweight, scalable, and interpretable metric for measuring the effectiveness of visual hand-to-robot transformations in preserving task semantics. Detailed results and further analyses are presented in Appendix D.
3.3 H2R Dataset Construction
We construct four large-scale robot-centric video datasets using the proposed H2R augmentation pipeline, each combining a specific robot embodiment with a human egocentric video dataset. These datasets are designed to support policy learning, visual representation learning, and embodiment generalization.
Applying H2R to SSv2. We begin with the SSv2 dataset, which contains 220,847 video clips of human actions with everyday objects, designed to help models understand fine-grained hand gestures. From the official training split, we select 62,500 videos that cover a wide variety of manipulation categories. For each video, we uniformly sample 16 keyframes, resulting in exactly 1,000,000 frames. Each sampled frame is passed through the H2R pipeline to replace the human hand with a simulated robotic arm. For the UR5 robot, we construct three distinct datasets with different end-effector configurations: H2R-UR5-SSv2-1M-Gripper uses a standard two-finger parallel gripper; H2R-UR5-SSv2-1M-Leaphand employs a four-finger anthropomorphic Leaphand; and H2R-UR5-SSv2-1M-Mix combines both embodiments at the frame level to increase visual diversity and generalization. For the Franka robot, which is configured only with a two-finger gripper, we construct H2R-Franka-SSv2-1M following the same augmentation strategy.
Applying H2R to Ego4D. To expand the applicability of our dataset to unconstrained, real-world human activity, we apply the same augmentation pipeline to the Ego4D dataset, which features long-form, first-person videos of natural daily behavior across diverse environments. Following the action-centric preprocessing strategy introduced by MPI zeng2024learningmanipulationpredictinginteraction , we extract 117,624 action clips from 2,486 Ego4D videos, each consisting of three keyframes. These are further processed to produce H2R-UR5-Ego4D-1M-Gripper, H2R-UR5-Ego4D-1M-Leaphand, and H2R-UR5-Ego4D-1M-Mix, using the same embodiment configuration logic as in the SSv2 case. For Franka, we similarly construct H2R-Franka-Ego4D-1M using the gripper configuration.
Each dataset contains approximately one million augmented images. In addition to the RGB images, we provide comprehensive frame-level metadata to support downstream use. Each entry includes: (1) the original label (from SSv2 or Ego4D), (2) the 3D hand keypoints, camera intrinsics, and extrinsics estimated by HaMeR, (3) the inpainted background image generated by LaMa, and (4) the final robot joint configurations and camera pose alignment used in the simulator for rendering. All robotic arms are rendered using the SAPIEN simulator, with camera viewpoint precisely matched to the original egocentric perspective using a two-step transformation between the HaMeR-estimated camera pose and the robot-world coordinate frame. Each dataset is released in a modular file structure, with paired directories for images, masks, and JSON metadata files. This organization ensures flexible access for both visual pre-training pipelines and embodied policy learning frameworks.
Figure 3: Examples of H2R Augmentation. Each column shows images before and after augmentation. The top row uses UR5 Leaphand, the middle row uses UR5 Gripper, and the bottom row uses Franka to replace human hands.
4 Simulation Experiment
4.1 Experiment Setup
Figure 4: Simulation Benchmark Overview. Visualization of the seven simulation tasks used for evaluation, including three from Robomimic (MoveCan, Square, Lift), one PushT task from Diffusion Policy, and three from RLBench (Close Box, Close Laptop Lid, Toilet Seat Down).
Encoder Pre-training. We adopt the MAE he2022mae ; xiao2022mvp and R3M nair2022r3m frameworks for pre-training, each employing a Vision Transformer (ViT) Base dosovitskiy2021imageworth16x16words model as the visual encoder. The training dataset used is either Ego4D (117K clips) or H2R-UR5-SSv2-1M-Mix (1M H2R-augmented images). Both MAE and R3M are pre-trained with 8$\times$×A800 GPUs. MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate. For policy training, we present the set of parameters that necessitate policy learning in Appendix  C.
Simulation Benchmark. For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning. Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic mandlekar2021matters , RLbench james2019rlbench , and Diffusion Policy chi2023diffusionpolicy . In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects. For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down. We also use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location. An overview of these simulation tasks is shown in Figure 4.
Table 1: Simulation Benchmark Results. Success rates (%) across imitation learning tasks in Robomimic, PushT, and RLBench environments, evaluated with MAE and R3M encoders before and after applying H2R. Blue represents an increase in task success rate, while Red represents a decrease. All subsequent tables follow the same rule.
Robomimic Diffusion Policy RLBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (SSv2) 54 25.5 94.5 58 59.2 0 10 0 3.3 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5 (+5%) 15 (+5.0%) 20 (+20.0%) 13.3 (+10%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5 (+5.0%) 20 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%)
4.2 Simulation Results
Table 1 shows that the encoders trained on H2R-UR5-SSv2-1M-Mix dataset significantly outperform those trained on the original Ego4D data across all simulation tasks. Specifically, for Robomimic tasks, the MAE encoder achieves an average success rate improvement of 10.2%, while the R3M encoder improves by 6.3%. In particular, the MoveCan task exhibits a substantial 25.5% increase for MAE trained on the H2R-1M dataset. For the PushT task, H2R leads to success rate gains of 5.3% (MAE) and 7.0% (R3M), further confirming the generalization benefits brought by our method. Moreover, in the RLBench benchmark, H2R consistently enhances performance, achieving an average improvement of 10.0% for MAE and 5.0% for R3M.
These results demonstrate that H2R-1M dataset delivers superior visual representations for imitation learning compared to large-scale human video datasets like Ego4D. H2R  achieve an improvement in the pre-training performance of the vision encoder on imitation learning by bridging the visual gap between human hands and robotic arms.
4.3 Performance on Other Datasets
Table 2: Results on Ego4D Dataset. Success rates (%) across imitation learning tasks in the PushT and RLBench environments.
Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D)) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.5 (-0.1%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
In addition to the H2R-UR5-SSv2-1M-Mix dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the H2R-UR5-Ego4D-1M-Mix dataset. The experimental results in the simulator are shown in Table 2, which denotes that H2R remains highly effective when applied to the Ego4D dataset.
5 Real World Experiment
Table 3: Real-world Task Results. We report the success rate (%) over real-world tasks for MAE and R3M.
Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 21.7 43.3 (+21.7%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%)
5.1 Experiment setup
Robot Setups. We validate the effectiveness of H2R in real-world manipulation tasks using a UR5 robotic arm with two different end effectors: Gripper RobotiqAdaptiveGrippers and Leaphand shaw2023leaphandlowcostefficient . Realsense IntelRealSenseL515 is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model. Real-world setups are shown in Figure 6.
Figure 5: Visualization of Real-world Manipulation Tasks. The left columns show Gripper tasks and the right columns show Leaphand tasks. Each task is illustrated with six frames, demonstrating the progression from the initial state to the completion of the manipulation.
Real-world Tasks. We set up six tasks for gripper manipulation and dexterous manipulation (Figure 5). For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed. For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened.
Data Collection. We collect expert demonstrations through human teleoperation. For gripper manipulation, we use keyboard-based teleoperation. For dexterous manipulation, we teleoperate Leaphand using a vision-based retargeting system. Due to differences in task complexity, robot embodiment, and teleoperation methods, we adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 4.
Table 4: Task-specific Data Collection and Evaluation Settings.
Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800
Policy Training. For policy training, we select the Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.1. DP and ACT are trained for 300 epochs for each gripper manipulation task and 3000 epochs for each dexterous manipulation task.
Evaluation. To ensure that the initial conditions are consistent throughout the evaluation, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstrations. During the evaluation, each real-world task is rolled out 20 times. We report the success rates of these tasks to assess model quality.
Gripper Setup
Leaphand Setup
Figure 6: Real-world Robot Setups. Experimental setups using a UR5 robot with two different end effectors: a parallel-jaw gripper (left) and a dexterous Leaphand hand (right).
5.2 Real-world Results
In the real-world tasks, we employ diffusion policy (DP) chi2023diffusionpolicy and ACT zhao2023act to drive robots in performing real-world tasks, where the visual encoders are pre-trained using MAE and R3M. Policy training details are presented in Appendix  C.
To evaluate the effectiveness and generalization of H2R, we conduct real-world experiments under two different robot embodiments: UR5 (same embodiment as downstream execution) / Franka (cross-embodiment generalization).
To assess the effectiveness of H2R, we use H2R-UR5-SSv2-1M-Mix during pre-training, which matches the embodiment used in policy training and evaluation. As shown in Table 3, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies. For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M), while for Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M). Similar trends are observed under ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), and Gripper tasks improving by 8.3% and 11.7%, respectively.
To evaluate cross-embodiment generalization, we use H2R-Franka-SSv2-1M dataset. Downstream policy training and evaluation are still conducted with UR5 Leaphand. As shown in Table 6, H2R still outperforms raw Ego4D despite the embodiment mismatch. Under DP, MAE improves from 40% to 43.3%, and R3M from 21.7% to 31.7%. Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%. Although Franka-based pretraining underperforms UR5, the results confirm H2R’s robustness to embodiment variations. Additional analysis, including failure case categorization and experiments on generalization under lighting perturbations, is provided in Appendix  E and  G.
5.3 Performance on Other Datasets
Table 5: Real-world Success Rates with Ego4D Pre-training. We report success rates (%) on real-world Leaphand tasks using visual encoders pre-trained on the Ego4D dataset. All downstream policies are trained using ACT.
Task MAE (Ego4D) MAE (H2R-Ego4D) R3M (Ego4D) R3M (H2R-Ego4D) Leaphand-GraspChicken 25 35 (+10%) 15 20 (+5%) Leaphand-StandCup 25 50 (+25%) 20 25 (+5%) Leaphand-OpenBox 35 45 (+10%) 30 40 (+10%) Average 28.3 43.3 (+15%) 21.7 28.3 (+6.7%)
We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented H2R-UR5-Ego4D-1M-Mix, and apply them to real-world Leaphand manipulation tasks using the ACT zhao2023act policy. As shown in Table 5, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M. These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
Table 6: Real-world Success Rates with H2R. We report the task success rates (%) using two robot embodiments (UR5, Franka) for data augmentation during pre-training. All models are evaluated on downstream tasks using the Leaphand end-effector, under MAE and R3M encoders with DP and ACT policies. Despite embodiment mismatch, H2R provides consistent gains over the baseline without augmentation.
Policy Task MAE (SSv2) MAE (H2R-UR5-SSv2) MAE (H2R-Franka-SSv2) R3M(SSv2) R3M (H2R-UR5-SSv2) R3M (H2R-Franka-SSv2) DP Leaphand-GraspChicken 40 55 (+15%) 35 (-5%) 10 35 (+25%) 20 (+10%) Leaphand-StandCup 35 60 (+25%) 50 (+15%) 20 50 (+30%) 30 (+10%) Leaphand-OpenBox 45 65 (+20%) 45 40 45 (+5%) 45 (+5%) Average 40 60 (+20%) 43.3 (+3.3%) 21.7 43.3 (+21.7%) 31.7 (+10%) ACT Leaphand-GraspChicken 45 50 (+5%) 50 (+5%) 10 35 (+25%) 40 (+30%) Leaphand-StandCup 25 50 (+25%) 50 (+25%) 20 60 (+40%) 25 (+5%) Leaphand-OpenBox 30 40 (+10%) 30 15 20 (+5%) 5 (-10%) Average 33.3 46.7 (+13.3%) 43.3 (+10%) 15 38.3 (+23.3%) 23.3 (+8.3%)
Table 7: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget). Task success rates (%) are reported under DP and ACT policies. Removing either component causes significant performance degradation, with red values showing the relative drop compared to full H2R.
Policy Task H2R H2R w/o Overlay H2R w/o Retarget DP Leaphand-GraspChicken 55 30 (-25%) 30 (-25%) Leaphand-StandCup 60 40 (-20%) 55 (-5%) Leaphand-OpenBox 65 20 (-45%) 45 (-20%) Average 60 30 (-30%) 43.3 (-16.7%) ACT Leaphand-GraspChicken 50 25 (-25%) 45 (-5%) Leaphand-StandCup 50 35 (-15%) 30 (-20%) Leaphand-OpenBox 40 25 (-15%) 15 (-25%) Average 46.7 28.3 (-18.3%) 30 (-16.7%)
5.4 Ablation study
To evaluate the effectiveness of each component in H2R, we conduct ablation studies on two time-consuming steps: (1) performing hand inpainting without overlaying a robotic arm (H2R w/o Overlay), and (2) overlaying the arm without precise alignment between the hand and the camera, instead using random pasting (H2R w/o Retarget). Table 7 shows the necessity and effectiveness of each component in H2R. The first step leads to a significant drop in success rate due to the loss of critical human-object interaction pixels after inpainting. The second step fails to provide accurate motion cues for the model and introduces visual mismatches with real-world manipulation tasks.
6 Conclusion
We propose H2R, a data augmentation technique that bridges the visual gap between human hand demonstrations and robotic arm manipulations by replacing human hands in first-person videos with robotic arm movements. Using 3D hand reconstruction and image inpainting models, H2R generates synthetic robotic arm manipulation sequences, making them more suitable for robot pre-training. Experiments across simulation benchmarks and real-world tasks demonstrate consistent improvements in success rates for encoders trained with various pre-training methods (e.g., MAE, R3M), highlighting the effectiveness and generalization of H2R. H2R enables efficient transfer of task knowledge from human demonstrations to robotic systems, reducing the reliance on costly robot-specific data collection. Finally, we expressed the broader impact of this work in Appendix H.
Appendix A Limitations
H2R is a simple data augmentation technique for robot pre-training from videos. There are several limitations, which open up possibilities for future work:
  1. 1.
    H2R augmentation pipeline is sensitive to the accuracy of pose estimation and hand segmentaion, which demands high-quality egocentric video datasets as prerequisites. This sensitivity becomes particularly pronounced in challenging scenarios involving low-light conditions or partial hand occlusions, where current estimation methods often fail to maintain required precision.
  2. 2.
    While our robotic platform incorporates diverse manipulator configurations spanning various robotic arm, grippers and dexterous hands, the current implementation remains limited to single-arm operation. Future work will expand this framework to dual-arm and humanoid robotics, significantly broadening the model’s application scope.
  3. 3.
    It would be valuable to extend our work on more tasks besides manipulation, like visual navigation and mobile manipulation.
Appendix B Details of Simulator Camera Position Alignment
We define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robot arm in the simulator. We build the coordinate system ${}^{W}\mathbf{I}_{H}$ based on the hand keypoints:
${}^{W}\mathbf{I}_{H}=\{^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}\}$ (2)
Where ${}^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}$ are unit vectors along the x-axes, y-axes and z-axes of the human hand coorinate system. With the keypoints get in HaMeR, we build the three axis of coordinates with the following functions:
${}^{w}\mathbf{i}_{H,x}$ $=^{w}\mathbf{i}_{0,9}$ (3)
${}^{w}\mathbf{i}_{H,y}$ $=^{w}\mathbf{i}_{0,9}\times^{w}\mathbf{i}_{0,13}$
${}^{w}\mathbf{i}_{H,z}$ $=^{w}\mathbf{i}_{H,x}\times^{w}\mathbf{i}_{H,y}$
Where ${}^{w}\mathbf{i}_{0,9}$ and ${}^{w}\mathbf{i}_{0,13}$ are unit vectors along the middle and ring fingers, respectively. In this notation, the first index (0) refers to the specific finger (middle or ring), and the second index (9 and 13) corresponds to the joint numbers along those fingers, as defined by the MANO model. Similarly, To construct the mapping from hand pose to robot arms, we need to get another coordinate system ${}^{W}\mathbf{I}_{S}$ in the simulator:
${}^{W}\mathbf{I}_{S}=\{^{w}\mathbf{i}_{S,x},^{w}\mathbf{i}_{S,y},^{w}\mathbf{i}_{S,z}\}$ (4)
The method of determining the axis of coordinates is the same:
${}^{w}\mathbf{i}_{S,x}$ $=^{w}\mathbf{i}_{0,2}$ (5)
${}^{w}\mathbf{i}_{S,y}$ $=^{w}\mathbf{i}_{0,2}\times^{w}\mathbf{i}_{0,3}$
${}^{w}\mathbf{i}_{s,z}$ $=^{w}\mathbf{i}_{S,x}\times^{w}\mathbf{i}_{S,y}$
Where $\mathbf{i}_{0,2},\mathbf{i}_{0,3}$ are unit vectors along robot fingers that correspond to human middle and ring fingers and the index corresponds to the joint numbers defined by MANO. We build the following two coordinate transformation matrix to construct the mapping:
${}^{W}_{H}\mathbf{R}$ $=\begin{pmatrix}{}^{W}\mathbf{I}_{H}&\mathbf{key}_{human}\\ \mathbf{O}&1\end{pmatrix}$ (6)
${}^{W}_{S}\mathbf{R}$ $=\begin{pmatrix}{{}^{W}\mathbf{I}_{S}}&{\mathbf{key}_{robot}}\\ \mathbf{O}&{1}\end{pmatrix}$
Where $\mathbf{key}_{human},\mathbf{key_{robot}}$ are the positions of human wrist and robot wrist. After obtaining the two coordinate systems, we need to determine the position of the camera in the simulator (${}^{W}\mathbf{cam}_{sim}$) and the position of the camera in the real world (${}^{H}\mathbf{cam}_{Real}$), thus we can ensure we get the same pose of the human hand and robot arms
${}^{H}\mathbf{cam}_{Real}$ $=^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (7)
${}^{S}\mathbf{cam}_{sim}$ $=^{H}\mathbf{cam}_{Real}$
${}^{W}\mathbf{cam}_{sim}$ $=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$
Figure 7: H2R samples. Visual comparison between original human data (top) and our augmented data (bottom).
Appendix C Policy Training Details
Policy Training in Simulation Experiment. For Robomimic tasks, we train for 200 steps and report the mean success rate. For tasks in RLBench, we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization. For the PushT task, we train the Diffusion Policy model for 200 epoches and report the success rate in the simulation environment. The training hyperparameters used in this work are identical to those described in the original paper.
Policy Training in Real-world Experiment. We select the Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section  4.1. DP and ACT are trained for 300 epochs for each gripper manipulation task and 3000 epochs for each dexterous manipulation task.
Appendix D Evaluation of H2R Effectiveness
In Figure 7 are six pairs of subfigures, each showing a human action and its corresponding augmented robotic action. The six action pairs in Figure 7 (left to right) are described as follows:
  • holding a yellow measuring tape at a construction site
  • holding a bottle
  • painting on the wall with a brush / Brushing paint on a wall
  • turning on the kitchen faucet
  • preparing painting materials on a table
The similarity scores computed by CLIP are presented in Table 8. Across these samples, the augmented images consistently exhibit higher similarity scores with the robot-centric action descriptions compared to the similarity scores between the unaugmented images and the human-centric action descriptions. This quantitative analysis demonstrates that the H2R augmentation effectively enhances the visual alignment between human and robotic actions.
Table 8: CLIP similarity scores. Higher values indicate better alignment between images and action descriptions.
Img1 Img2 Img3 Img4 Img5 Img6
ori 30.6 23.7 31.2 31.7 27.5 28.7
aug 32.6 28.9 32.3 32.0 28.7 31.0
Stack
Stand Cup
Figure 8: Failure case visualizations: Stack and Stand Cup. We visualize real-world manipulation executions for two downstream tasks: Stack (top) and Stand Cup (bottom). These images provide qualitative insights into the performance and failure modes of the policy in real deployment, highlighting challenges such as object misalignment, perception noise, and grasp precision.
Appendix E Failure Case Analysis
Table 9: Sub-goal.Task-specific sub-goal evaluation. To gain fine-grained insights into policy performance, we design a manual rubric covering key sub-goals for each manipulation task. Each cell reports the number of successful vs. unsuccessful attempts (Y/N) over 20 evaluation trials. Results show that models enhanced with H2R consistently accomplish more sub-goals across tasks compared to their baseline counterparts, demonstrating improved robustness in real-world execution. Bold numbers indicate better performance between paired models.
Task Sub-goal MAE(Y/N) MAE+H2R(Y/N) R3M(Y/N) R3M+H2R(Y/N) Gripper-PickCube Overall success? 9/11 13/7 8/12 10/10 Pick up the cube? 14/6 15/5 11/9 13/7 Gripper-Stack Overall success? 10/10 11/9 11/9 14/6 Pick up the cube? 13/7 16/4 13/7 17/3 Gripper-CloseBox Overall success? 11/9 10/10 9/11 13/7 Place the cube in the bow? 12/8 14/6 12/8 15/5 Pick up the cube? 14/6 14/6 12/8 15/5 Leaphand-GraspChicken Overall success? 8/12 11/9 2/18 7/13 Pick up the chicken? 13/7 14/6 3/17 10/10 Leaphand-StandCup Overall success? 7/13 12/8 4/16 10/10 Pick up the cup? 12/8 18/2 12/8 15/5 Leaphand-OpenBox Overall success? 9/11 13/7 8/12 9/11 Identify contact location? 14/6 16/4 10/10 10/10
To better understand the limitations of our policy and the challenges encountered in real-world deployments, we present a qualitative analysis of failure cases from two representative tasks: a gripper-based task (Gripper-Stack) and a dexterous manipulation task (Gripper-StandCup). Figure 8 illustrates typical failure modes observed during execution.
In the Gripper-Stack task, we identify two major failure scenarios:
Case I: Grasp Failure Unnoticed. The robot arm fails to successfully grasp the blue cube. However, the policy proceeds as if the object had been grasped, moving toward the yellow cube and attempting to perform the stacking operation. This leads to a complete task failure.
Case II: Misaligned Placement. The robot successfully grasps the blue cube but fails to align it correctly on top of the yellow cube during the stacking phase, resulting in an unstable or failed placement.
In the Stand Cup task, similar issues emerge due to perception and control limitations:
Case I: Grasp Position Error. The Leaphand end-effector attempts to grasp the cup but fails to target the correct contact region. As a result, the cup slips out of the grasp during lifting, preventing task completion.
Case II: Insufficient Lifting Trajectory. Even when the grasp is successful, the lifting motion lacks sufficient amplitude or stability to fully stand the cup upright. The cup either tips over or fails to stand securely.
To enable fine-grained evaluation of policy performance and gain deeper insights into failure cases, we designed a task-specific evaluation rubric. Table 9 displays our rubric that the evaluator filled out when rolling out different policies. Take the DP policy as an example, the results in Table 9 demonstrate that H2R-augmented visual representation models not only improve overall success rates in real-world tasks, but also allow to accomplish more than half of the task consistently.
Table 10: Robomimic Experiment result. We report the success rate (%) over IL-based tasks for MAE and R3M Robomimic.
MoveCan Square Lift Average PushT MAE 54 25.5 94.5 58 59.2 MAE+CutMix1 72.0 (+18.0%) 30.0 (+4.5%) 95.0 (+0.5%) 65.7 (+7.7%) 37.5 (-21.7%) MAE+CutMix2 58.0 (+4.0%) 36.0 (+10.5%) 90.0 (-4.5%) 61.3 (+3.3%) 40.0 (-19.2%) MAE+CutMix3 78.0 (+24.0%) 32.0 (+9.3%) 92.0 (-2.5%) 67.3 (+2.7%) 42.0 (-17.2%) MAE+H2R 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) R3M 59.5 20.5 85 55 15 R3M+CutMix1 69.5 (+10.0%) 30.0 (+9.5%) 91.0 (+6.0%) 63.5 (+8.5%) 19.0 (+4.0%) R3M+CutMix2 66.0 (+6.5%) 26.0 (+5.5%) 83.0 (-2.0%) 58.3 (+3.3%) 17.0 (+2.0%) R3M+CutMix3 68.0 (+8.5%) 26.0 (+5.5%) 84.0 (-1.0%) 59.3 (+4.3%) 14.0 (-1.0%) R3M+H2R 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%)
Appendix F Additional Ablation Study in Simulation Experiment
In addition to pre-training on the H2R data and raw data, we also applied a simple CutMix baseline to demonstrate the effectiveness of using the robotic arm to cover the human hand, which overlays a fixed set of specific images of robotic arms with grippers onto the original images, ensuring that the overlaid images cover the human hands as much as possible, without exceeding the detected bounding box. Our H2R is different from such baseline by employing robot hand construction to better match the pose of the hand and arm in the images. Based on the type of robotic arm used in CutMix, we categorize the augmented set into three types: CutMix1 represents the UR5 robotic arm, CutMix2 refers to the Franka robotic arm, and CutMix3 combines both the UR5 and Franka robotic arms.
From Table 10, we observe that the encoder trained on H2R processed data shows consistent improvements across various tasks compared to the encoder trained on the original data, with the average success rate on all tasks ranging from 0.9% to 10.2%. Especially for the more challenging MoveCan task, it can improve the success rate by 25.5%. Additionally, while encoders trained on the relatively simple CutMix data show improvement on tasks in Robomimic, their performance in the PushT task remains slightly worse than the encoders trained on original data. These results demonstrate the effectiveness of using the robotic arm to cover the human hand in video data, as well as the effectiveness of H2R in imitation learning.
Table 11: Generalization under Lighting Variations. Success rates (%) under real-world lighting disturbances. Models trained with additional lighting augmentations (LightAug) show significant improvements in robustness compared to baseline and H2R-only models.
Tasks MAE MAE+H2R MAE+H2R+LightAug R3M R3M+H2R R3M+H2R+LightAug Leaphand-GraspChicken 10 10 20 0 0 10 Leaphand-StandCup 20 25 40 5 15 20 Leaphand-OpenBox 5 0 25 10 10 10 Average 11.7 11.7 28.3 5 8.3 13.3
Appendix G Generalization on Light Condition.
Figure 9: Lighting Setup in Real-World Experiments. Light configuration used in real-world evaluations.
To evaluate the generalization under varying lighting conditions, we introduce illumination disturbances during evaluation, as illustrated in Figure 9. Additionally, during training, we incorporate randomized lighting with varying directions and colors into the simulation environment for data augmentation. We compare three settings: no augmentation(MAE, R3M), H2R augmentation(MAE+H2R, R3M+H2R), and H2R with lighting disturbances(MAE+H2R+LightAug, R3M+H2R+LightAug). As shown in Table 11, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.
Appendix H Broader Impact
In this work, we proposed a a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric video. We focused on the theme above and doesn’t have direct social impact. We hope that through our data augmentation methods, more egocentric human datasets can be used for higher-level robot manipulation models, thereby promoting progress in the field of embodied intelligence.

### v3

@@
 We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data.
-Based on this, we construct and release a family of 1M-scale datasets covering multiple robot embodiments (UR5 with gripper/Leaphand, Franka) and data sources (SSv2, Ego4D).
 To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions.
-We validate H2R across three simulation benchmarks—Robomimic, RLBench, and PushT and real-world manipulation tasks with a UR5 robot equipped with Gripper and Leaphand end-effectors.
-H2R consistently improves downstream success rates, yielding gains of 5.0%–10.2% in simulation and 6.7%–23.3% in real-world tasks across various visual encoders and policy learning methods.
+We validate H2R across four simulation benchmarks—Robomimic, RLBench, PushT and CortexBench, and real-world manipulation tasks with UR5 and Dual-Arm Franka/UR5.
+H2R consistently improves downstream success rates, yielding gains of 1.3%–10.2% in simulation and 3.3%–23.3% in real-world tasks across various visual encoders and policy learning methods.
+We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with VLA models.
 These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains.
-Figure 1: Overview of H2R.
-H2R is a data augmentation technique designed to enhance robot pre-training by converting first-person human hand operation videos into robot-centric visual data.
-By bridging the visual domain gap, H2R improves pre-trained visual encoders for downstream robot policies (imitation learning), validated across simulation benchmarks and real-world robotic tasks.
+The project page is available in https://sites.google.com/view/h2r-robotics.
 1 Introduction
 Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics.
-Inspired by the remarkable accomplishments of large scale pre-training in computer vision he2022mae ; nair2022r3m ; zeng2024learningmanipulationpredictinginteraction ; karamcheti2023voltron ; datacomp and natural language processing devlin2018bert ; yu2022pointbert ; openai2024gpt4technicalreport ; anil2023palm2technicalreport ; li2024datacomplm , many efforts have been devoted to harness large-scale data to construct generalizable representations in the robotics field embodimentcollaboration2024openxembodimentroboticlearning ; brohan2023rt1roboticstransformerrealworld ; brohan2023rt2visionlanguageactionmodelstransfer .
-Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive duan2023ar2 ; jang2022bczzeroshottaskgeneralization ; brohan2023rt1roboticstransformerrealworld ; khazatsky2024droidlargescaleinthewildrobot ; fang2023rh20tcomprehensiveroboticdataset ; gao2024efficientdatacollectionrobotic ; lee2021pickandplacetacklingroboticstacking ; herzog2023deeprlscalesorting ; kalashnikov2021mtoptcontinuousmultitaskrobotic ; duan2024manipulate ; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
-Recent works nair2022r3m ; xiao2022mvp ; zeng2024learning analyze such egocentric human video datasets such as Ego4D grauman2022ego4d , SSv2 goyal2017something , and Epic Kitchens damen2018scaling with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations.
+Inspired by the remarkable accomplishments of large scale pre-training in computer vision [he2022mae, nair2022r3m, zeng2024learningmanipulationpredictinginteraction, karamcheti2023voltron, datacomp] and natural language processing [devlin2018bert, yu2022pointbert, openai2024gpt4technicalreport, anil2023palm2technicalreport, li2024datacomplm], many efforts have been devoted to harnessing large-scale data to construct generalizable representations in the robotics field [embodimentcollaboration2024openxembodimentroboticlearning, brohan2023rt1roboticstransformerrealworld, brohan2023rt2visionlanguageactionmodelstransfer].
+Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive [duan2023ar2, jang2022bczzeroshottaskgeneralization, brohan2023rt1roboticstransformerrealworld, khazatsky2024droidlargescaleinthewildrobot, fang2023rh20tcomprehensiveroboticdataset, gao2024efficientdatacollectionrobotic, lee2021pickandplacetacklingroboticstacking, herzog2023deeprlscalesorting, kalashnikov2021mtoptcontinuousmultitaskrobotic, duan2024manipulate]; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
+Recent works [nair2022r3m, xiao2022mvp, zeng2024learning] analyze such egocentric human video datasets such as Ego4D [grauman2022ego4d], SSv2 [goyal2017something], and Epic Kitchens [damen2018scaling] with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations.
 However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems.
@@
 This would result in increased complexity during the fine-tuning process and suboptimal performance.
-To mitigate this issue, we propose H2R (as shown in Figure 1), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation.
+To mitigate this issue, we propose H2R (as shown in Figure LABEL:fig:pipeline), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation.
 H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video.
-Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR pavlakos2023reconstructinghands3dtransformers to accurately detect the position and posture of the human hand in egocentric videos.
+Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [pavlakos2023reconstructinghands3dtransformers] to accurately detect the position and posture of the human hand in egocentric videos.
 Then, we simulate the same robot state in simulators to obtain the mask of robot actions.
-In the second stage, we use the Segment Anything Model kirillov2023segment to automatically separate human hand from background, and use the inpainting model LaMa suvorov2021resolutionrobustlargemaskinpainting to fill the removed hand mask.
+In the second stage, we use the Segment Anything Model [kirillov2023segment] to automatically separate human hand from background, and use the inpainting model LaMa [suvorov2021resolutionrobustlargemaskinpainting] to fill the removed hand mask.
 After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video.
@@
 It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
-Based on this pipeline, we construct and release a series of large-scale robot-centric datasets (H2R-1M), each containing approximately 1 million augmented frames.
-These datasets cover multiple robot embodiments (UR5 with gripper/Leaphand, Franka) and human video sources (SSv2, Ego4D), and provide modular annotations to support downstream robotic learning.
 To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics.
@@
 These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice.
-For the pre-training stage, we apply MAE he2022mae and R3M nair2022r3m frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video.
+For the pre-training stage, we apply MAE [he2022mae] and R3M [nair2022r3m] frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video.
 The pretrained visual representations are used in downstream imitation learning pipelines.
 We conduct comprehensive evaluations across both simulation and real-world manipulation tasks.
-In simulation, we evaluate pre-trained visual encoders on three benchmark suites: Robomimic, RLBench, and PushT.
-H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, and +5.3% for MAE and +7.0% for R3M on the PushT task.
-In real-world experiments, we deploy H2R-enhanced encoders on a UR5 robot equipped with both Gripper and Leaphand end-effectors, evaluated over six manipulation tasks under two policy frameworks: Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act .
-H2R brings consistent improvements, with average success rate gains of +13.3% (MAE) and +17.0% (R3M) under DP, and +10.8% (MAE) and +17.5% (R3M) under ACT.
+In simulation, we evaluate pre-trained visual encoders on four benchmark suites: Robomimic, RLBench, PushT and CortexBench.
+H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, +5.3% for MAE and +7.0% for R3M on the PushT task, and +5.3% for MAE and +1.3% for R3M on the CortexBench task.
+In real-world experiments, we deploy H2R-enhanced encoders on different robot and end-effector settings, evaluated over nine manipulation tasks under two policy frameworks: Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act].
+H2R brings consistent improvements, with average success rate gains of +10.0% (MAE) and +14.1% (R3M) under DP, and +12.8% (MAE) and +13.3% (R3M) under ACT.
 H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively.
-These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulated and real-world robotic manipulation settings.
-Our paper provides three contributions:
-  • We propose a data-centric pipeline, H2R, to mitigate the gap between human and robot hands when utilizing large-scale egocentric video datasets to pre-train generalizable visual features for robots.
-  • We construct and release diverse robot-centric datasets (H2R-1M), combining multiple embodiments and egocentric video sources to support robot-compatible visual pretraining.
-  • We demonstrate the effectiveness of H2R through extensive experiments on closed-loop benchmarks.
+Beyond these evaluations, we also test H2R’s ability to support cross-embodiment transfer by pre-training encoders on Franka data and deploying them on UR robots, indicating that H2R improves representation transfer across different morphologies.
+In addition, fine-tuning VLA backbones on H2R-generated datasets yields markedly higher success rates on dual-arm UR tasks than fine-tuning on human-hand videos or using no fine-tuning.
+These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulation and real-world manipulation settings.
 2 Related Work
 Robot Imitation Learning.
-Data-driven policy learning sergey2018learning ; luo2024hilserl ; xiao2022mvp ; yang2024equibotsim3equivariantdiffusionpolicy ; chi2023diffusionpolicy ; embodimentcollaboration2024openxembodimentroboticlearning has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation.
-Imitation learning chi2023diffusionpolicy ; ze2024dp3 ; zhao2023act ; mandlekar2021matters trains policies from successful demonstrations, often supervised by behavior cloning torabi2018behavioral ; florence2021implicit objectives.
-ACT zhao2023act addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models chi2023diffusionpolicy ; ze2024dp3 are introduced to handle the inherent multimodality of robot motions.
-In addition, CordViP fu2025cordvip leverages 3D object-robot correspondences to enhance dexterous manipulation.
+Data-driven policy learning [sergey2018learning, luo2024hilserl, xiao2022mvp, yang2024equibotsim3equivariantdiffusionpolicy, chi2023diffusionpolicy, embodimentcollaboration2024openxembodimentroboticlearning] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation.
+Imitation learning [chi2023diffusionpolicy, ze2024dp3, zhao2023act, mandlekar2021matters] trains policies from successful demonstrations, often supervised by behavior cloning [torabi2018behavioral, florence2021implicit] objectives.
+ACT [zhao2023act] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [chi2023diffusionpolicy, ze2024dp3] are introduced to handle the inherent multimodality of robot motions.
+In addition, CordViP [fu2025cordvip] leverages 3D object-robot correspondences to enhance dexterous manipulation.
 Visual Encoder Pretraining for Robotics.
 Visual pretraining improves generalization of robotic policies across diverse tasks.
-Researchers have explored architectural designs He_2016_CVPR ; dosovitskiy12021an , training objectives He_2022_CVPR ; chen2020simple ; chen2020big , and dataset compositions deng2009imagenet ; lin2014microsoft ; shao2019object365 ; Schuhmann2021LAION400MOD .
-PVR-Control parisi22a shows that pretrained visual representations can outperform direct state-based policies.
-RPT radosavovic2023robot tokenizes observations to enable masked prediction pretraining.
-Methods like MVP radosavovic2022realworld and R3M nair2022r3m utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning.
-Voltron karamcheti2023voltron demonstrates the use of MAE and contrastive learning for hierarchical robot control.
+Researchers have explored architectural designs [He_2016_CVPR, dosovitskiy12021an], training objectives [He_2022_CVPR, chen2020simple, chen2020big], and dataset compositions [deng2009imagenet, lin2014microsoft, shao2019object365, Schuhmann2021LAION400MOD].
+PVR-Control [parisi22a] shows that pretrained visual representations can outperform direct state-based policies.
+RPT [radosavovic2023robot] tokenizes observations to enable masked prediction pretraining.
+Methods like MVP [radosavovic2022realworld] and R3M [nair2022r3m] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning.
+Voltron [karamcheti2023voltron] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
 Cross-Domain Visual Alignment.
 Bridging the domain gap between human and robot visual inputs remains a major challenge.
-WHIRL bahl2022humantorobotimitationwild matches task structure from third-person views, while RoVi-Aug chen2024roviaugrobotviewpointaugmentation and Mirage chen2024mirage manipulate appearance via segmentation or image-space preprocessing.
-EgoMimic kareer2024egomimicscalingimitationlearning removes hands and normalizes views to align egocentric perspectives.
+WHIRL [bahl2022humantorobotimitationwild] matches task structure from third-person views, while RoVi-Aug [chen2024roviaugrobotviewpointaugmentation] and Mirage [chen2024mirage] manipulate appearance via segmentation or image-space preprocessing.
+EgoMimic [kareer2024egomimicscalingimitationlearning] removes hands and normalizes views to align egocentric perspectives.
 3 H2R: Human-to-Robot Data Augmentation
@@
 This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains.
+We first examine the position and pose of the hand and arm in the image, then map this pose to the robotic arm and end-effector in the simulator.
+Next, we erase the original hand and arm from the image and overlay a rendered image of the simulated robot performing the same pose onto the original image in place of the removed hand and arm.
 Figure 3 illustrates examples of human hand videos that have been processed using H2R.
-These examples include different types of robots (UR5 and Franka) equipped with a variety of end-effectors, such as dexterous hands and grippers.
-We further introduce a CLIP-based metric to evaluate semantic consistency, and construct a family of large-scale datasets that combine different robot embodiments (UR5 with gripper, Leaphand, or both; Franka with gripper) and egocentric sources (SSv2 and Ego4D).
 3.1 H2R Data Augmentation Pipeline
@@
 In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information.
-We adopt HaMeR pavlakos2023reconstructinghands3dtransformers , a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal.
+We adopt HaMeR [pavlakos2023reconstructinghands3dtransformers], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal.
 Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
-Human Arm and Hand Removal.
-After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background.
-We use Segment Anything Model (SAM) kirillov2023segment to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR.
-Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa suvorov2021resolutionrobustlargemaskinpainting , a powerful inpainting model, to fill in the removed hand-arm region.
-This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
 Robotic Arm and End Effector Construction.
-This step involves constructing the robotic arm and different end effectors.
-There are two common types of end-effectors in robotic manipulation: grippers and dexterous hands.
-In our pipeline, we handle them differently based on their structural characteristics.
-For dexterous hands, which have multiple degrees of freedom and resemble human hands, we compute the joint angles by analyzing the hand keypoints predicted by HaMeR.
-Specifically, each finger joint angle is determined by the angle formed between three consecutive keypoints corresponding to that finger segment.
-For grippers, which typically consist of two or more parallel fingers, we determine the degree of opening or closing based on the Euclidean distance between the relevant fingertips in the hand keypoints.
-However, since hand keypoints alone are insufficient to fully define the robotic arm’s complete configuration, especially for joints that do not directly correspond to hand movement, we manually assign plausible values to these undetermined joint positions to construct a reasonable arm pose.
+This step involves building the robotic arm and its end-effectors, which are typically either grippers or dexterous hands.
+For dexterous hands, we estimate joint angles from hand keypoints predicted by HaMeR.
+Each finger joint angle is calculated using three consecutive keypoints along that finger.
+For grippers, we determine how open or closed they are based on the Euclidean distance between the corresponding fingertips.
+Since hand keypoints don’t capture the full arm pose (especially for joints unrelated to the hand), we manually set reasonable values for the remaining arm joints to complete a plausible robot configuration.
 Simulator Camera Position Alignment.
@@
 $^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (1)
-A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix  B.
+A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix.
 Robot Hand Rendering and Copy-paste.
-After setting the camera, the segmentation mask of the robotic arm is obtained by shooting with the camera.
+After setting the camera, the segmentation mask of the robotic arm is obtained by rendering with the simulator camera.
 We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image.
@@
 By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
+Human Arm and Hand Removal.
+After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background.
+We use Segment Anything Model (SAM) [kirillov2023segment] to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR.
+Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [suvorov2021resolutionrobustlargemaskinpainting], a powerful inpainting model, to fill in the removed hand-arm region.
+This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
 3.2 Data Quality Evaluation
-To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP radford2021learningtransferablevisualmodels .
+Figure 2: H2R samples. Visual comparison between original human data (top) and our augmented data (bottom).
+To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP [radford2021learningtransferablevisualmodels].
 This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks.
 For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective.
-The human-centric template is defined as: A human is [action]”; the robot-centric template is defined as: A robotic arm is [action]” .
-Here, [action] is a natural language phrase describing the high-level behavior, such as holding a bottle” or “brushing paint on a wall.”
-We use the CLIP ViT-B/32 model to compute cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt.
+The human-centric template is defined as: A human is [action]; the robot-centric template is defined as: A robotic arm is [action].
+Here, [action] is a natural language phrase describing the high-level behavior, such as ”holding a bottle” or ”brushing paint on a wall”.
+We use the CLIP ViT-B/32 model to compute the cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt.
 This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description.
@@
 All textual prompts are manually curated to maintain consistency across frames and embodiments.
-This evaluation protocol provides a lightweight, scalable, and interpretable metric for measuring the effectiveness of visual hand-to-robot transformations in preserving task semantics.
-Detailed results and further analyses are presented in Appendix D.
-3.3 H2R Dataset Construction
-We construct four large-scale robot-centric video datasets using the proposed H2R augmentation pipeline, each combining a specific robot embodiment with a human egocentric video dataset.
-These datasets are designed to support policy learning, visual representation learning, and embodiment generalization.
-Applying H2R to SSv2.
-We begin with the SSv2 dataset, which contains 220,847 video clips of human actions with everyday objects, designed to help models understand fine-grained hand gestures.
-From the official training split, we select 62,500 videos that cover a wide variety of manipulation categories.
-For each video, we uniformly sample 16 keyframes, resulting in exactly 1,000,000 frames.
-Each sampled frame is passed through the H2R pipeline to replace the human hand with a simulated robotic arm.
-For the UR5 robot, we construct three distinct datasets with different end-effector configurations: H2R-UR5-SSv2-1M-Gripper uses a standard two-finger parallel gripper; H2R-UR5-SSv2-1M-Leaphand employs a four-finger anthropomorphic Leaphand; and H2R-UR5-SSv2-1M-Mix combines both embodiments at the frame level to increase visual diversity and generalization.
-For the Franka robot, which is configured only with a two-finger gripper, we construct H2R-Franka-SSv2-1M following the same augmentation strategy.
-Applying H2R to Ego4D.
-To expand the applicability of our dataset to unconstrained, real-world human activity, we apply the same augmentation pipeline to the Ego4D dataset, which features long-form, first-person videos of natural daily behavior across diverse environments.
-Following the action-centric preprocessing strategy introduced by MPI zeng2024learningmanipulationpredictinginteraction , we extract 117,624 action clips from 2,486 Ego4D videos, each consisting of three keyframes.
-These are further processed to produce H2R-UR5-Ego4D-1M-Gripper, H2R-UR5-Ego4D-1M-Leaphand, and H2R-UR5-Ego4D-1M-Mix, using the same embodiment configuration logic as in the SSv2 case.
-For Franka, we similarly construct H2R-Franka-Ego4D-1M using the gripper configuration.
-Each dataset contains approximately one million augmented images.
-In addition to the RGB images, we provide comprehensive frame-level metadata to support downstream use.
-Each entry includes: (1) the original label (from SSv2 or Ego4D), (2) the 3D hand keypoints, camera intrinsics, and extrinsics estimated by HaMeR, (3) the inpainted background image generated by LaMa, and (4) the final robot joint configurations and camera pose alignment used in the simulator for rendering.
-All robotic arms are rendered using the SAPIEN simulator, with camera viewpoint precisely matched to the original egocentric perspective using a two-step transformation between the HaMeR-estimated camera pose and the robot-world coordinate frame.
-Each dataset is released in a modular file structure, with paired directories for images, masks, and JSON metadata files.
-This organization ensures flexible access for both visual pre-training pipelines and embodied policy learning frameworks.
+Figure 2 contains six pairs of subfigures, each showing a human action and its corresponding augmented robotic action.
+The six action pairs in Figure 2 (left to right) are described as follows: (1) Holding a yellow measuring tape at a construction site; (2) Holding a bottle; (3) Painting on the wall with a brush; (4) Rushing paint on a wall; (5) Turning on the kitchen faucet; (6) Preparing painting materials on a table.
+The CLIP similarity scores are reported in Table 1.
+Across all samples, H2R-augmented images consistently achieve higher similarity with robot-centric action descriptions than original images do with human-centric ones.
+This confirms that H2R augmentation effectively improves visual-semantic alignment between human demonstrations and robotic execution.
+Table 1: CLIP similarity scores. Higher values indicate better alignment between images and action descriptions.
+Img-1 Img-2 Img-3 Img-4 Img-5 Img-6
+Ori 30.6 23.7 31.2 31.7 27.5 28.7
+Aug 32.6 28.9 32.3 32.0 28.7 31.0
+This evaluation provides a lightweight, scalable, and interpretable metric for assessing how well visual hand-to-robot transformations preserve task semantics.
+To quantify data quality, we sample 1,000 image pairs (original and H2R-augmented).
+For each original image, we use Qwen-2.5-VL [bai2025qwen25vltechnicalreport] to generate a verb-noun action description (e.g., “washing bowl”).
+We then form two prompts: “A human is [action]” and “A robot arm is [action]”.
+Using CLIP ViT-B/32, we compute the cosine similarity between: (1) the human prompt and the original image, and (2) the robot prompt and the H2R-augmented image.
+We report the average similarity over all 1,000 samples.
+The raw images achieve an average score of 28.01, while H2R-augmented images reach 29.83.
+This demonstrates that H2R-augmented images exhibit stronger semantic alignment with robot-centric action descriptions, confirming improved visual-semantic coherence for robotic manipulation.
+The results validate our data curation pipeline as effective for generating high-quality, robot-relevant visual data at scale.
 Figure 3: Examples of H2R Augmentation.
@@
 4 Simulation Experiment
-4.1 Experiment Setup
-Figure 4: Simulation Benchmark Overview.
-Visualization of the seven simulation tasks used for evaluation, including three from Robomimic (MoveCan, Square, Lift), one PushT task from Diffusion Policy, and three from RLBench (Close Box, Close Laptop Lid, Toilet Seat Down).
 Encoder Pre-training.
-We adopt the MAE he2022mae ; xiao2022mvp and R3M nair2022r3m frameworks for pre-training, each employing a Vision Transformer (ViT) Base dosovitskiy2021imageworth16x16words model as the visual encoder.
-The training dataset used is either Ego4D (117K clips) or H2R-UR5-SSv2-1M-Mix (1M H2R-augmented images).
-Both MAE and R3M are pre-trained with 8$\times$×A800 GPUs.
+We adopt the MAE [he2022mae, xiao2022mvp] and R3M [nair2022r3m] frameworks for pre-training, each employing a Vision Transformer (ViT) Base [dosovitskiy2021imageworth16x16words] model as the visual encoder.
+The training dataset used is a subset of Ego4D (117K clips) or SSv2 (1M images).
+Both MAE and R3M are pre-trained with $\timesA8$00 GPUs.
 MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate.
-For policy training, we present the set of parameters that necessitate policy learning in Appendix  C.
+Policy Training.
+We train for 200 steps and report the mean success rate for Robomimic tasks.
+For tasks in RLBench [james2019rlbench], we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization.
+For the PushT task, we train the Diffusion Policy model for 200 epochs and report the success rate in the simulation environment.
+For VC-1 Cortex benchmark, we train policies using behavior cloning composed of a 3-layer MLP for 100 epochs.
+The training hyperparameters used in this work are identical to those described in the original paper.
 Simulation Benchmark.
 For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning.
-Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic mandlekar2021matters , RLbench james2019rlbench , and Diffusion Policy chi2023diffusionpolicy .
+Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic [mandlekar2021matters], RLbench [james2019rlbench], and Diffusion Policy [chi2023diffusionpolicy].
 In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects.
 For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down.
-We also use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location.
-An overview of these simulation tasks is shown in Figure 4.
+We use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location.
+We also use three MetaWorld tasks: Assembly, ButtonPress and Hammer from VC-1 Cortex benchmark for evaluation.
+An overview of these simulation tasks is shown in Appendix.
 Table 2: Simulation Benchmark Results.
-Success rates (%) across imitation learning tasks in Robomimic, PushT, and RLBench environments, evaluated with MAE and R3M encoders before and after applying H2R.
+Success rates (%) across imitation learning tasks in Robomimic, PushT, RLBench and CortexBench, evaluated with MAE and R3M encoders before and after applying H2R.
 Blue represents an increase in task success rate, while Red represents a decrease.
 All subsequent tables follow the same rule.
-Robomimic Diffusion Policy RLBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (SSv2) 54 25.5 94.5 58 59.2 0 10 0 3.3 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5 (+5%) 15 (+5.0%) 20 (+20.0%) 13.3 (+10%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5 (+5.0%) 20 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%)
+Robomimic Diffusion Policy RLBench CortexBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average Assembly ButtonPress Hammer Average MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) 88.0(+4.0%) 88.0(+8.0%) 100.0 (+4.0%) 92.0(+5.3%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 76.0 56.0 88 73.3 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5.0 (+5.0%) 20.0 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%) 68.0 (-8.0%) 60.0 (+4.0%) 96.0 (+8.0%) 74.7 (+1.3%)
 4.1 Simulation Results
-Table 1 shows that the encoders trained on H2R-UR5-SSv2-1M-Mix dataset significantly outperform those trained on the original Ego4D data across all simulation tasks.
-Specifically, for Robomimic tasks, the MAE encoder achieves an average success rate improvement of 10.2%, while the R3M encoder improves by 6.3%.
-In particular, the MoveCan task exhibits a substantial 25.5% increase for MAE trained on the H2R-1M dataset.
-For the PushT task, H2R leads to success rate gains of 5.3% (MAE) and 7.0% (R3M), further confirming the generalization benefits brought by our method.
-Moreover, in the RLBench benchmark, H2R consistently enhances performance, achieving an average improvement of 10.0% for MAE and 5.0% for R3M.
-These results demonstrate that H2R-1M dataset delivers superior visual representations for imitation learning compared to large-scale human video datasets like Ego4D.
-H2R  achieve an improvement in the pre-training performance of the vision encoder on imitation learning by bridging the visual gap between human hands and robotic arms.
-4.3 Performance on Other Datasets
+Table 2 shows encoders trained on augmented data H2R-SSv2 significantly outperform those on original SSv2 across all simulation tasks.
+For Robomimic tasks, MAE achieves a 10.2% average success rate improvement and R3M 6.3%, with MAE seeing a substantial 25.5% gain on MoveCan.
+On PushT, H2R boosts MAE by 5.3% and R3M by 7.0%.
+In RLBench, MAE improves its performance by 10.0% on average and R3M by 5.0%.
+In CortexBench, H2R improves MAE by 5.3% and R3M by 1.3%.
+These results confirm that the H2R-augmented dataset delivers superior visual representations for imitation learning to those learned from large-scale human video datasets like SSv2, as H2R  bridges the visual gap between human hands and robotic arms to enhance vision encoder pre-training performance.
 Table 3: Results on Ego4D Dataset. Success rates (%) across imitation learning tasks in the PushT and RLBench environments.
-Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D)) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.5 (-0.1%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
-In addition to the H2R-UR5-SSv2-1M-Mix dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the H2R-UR5-Ego4D-1M-Mix dataset.
-The experimental results in the simulator are shown in Table 2, which denotes that H2R remains highly effective when applied to the Ego4D dataset.
+Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.6 (+0.0%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
+In addition to SSv2 dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the Ego4D dataset.
+The experimental results in the simulator are shown in Table 3, which indicates that H2R remains highly effective when applied to the Ego4D dataset.
+4.2 Pretraining Model on Robotic Datasets
+Table 4: Comparison with model pretrained on robotic datasets.
+All visual encoder weights are initialized using the R3M pretraining method, with the “R3M-H2R” variant pretrained on the H2R-augmented SSv2 dataset.
+Robomimic MoveCan Square Lift Average R3M 59.5 20.5 85.0 55.0 R3M-H2R 61.5 37.5 85.0 61.3 R3M-DROID 54.0 22.0 96.0 56.7
+To further evaluate how H2R compares to models pretrained directly on robotic datasets, we refer to a recent study [jiang2024robotspretrainrobotsmanipulationcentric], which conducted a systematic comparison between visual representations pretrained on large-scale egocentric human video datasets and those trained on robotic datasets such as DROID [khazatsky2024droidlargescaleinthewildrobot].
+The experimental results, summarized in Table 4, show that models pretrained on robotic datasets often provide limited or even degraded performance on downstream manipulation Robomimic benchmarks.
+Moreover, despite being smaller and simulated rather than real-robot data, the H2R dataset outperforms the DROID dataset in pre-training effectiveness, indicating its strong potential for robotic model pre-training and reflecting its high data quality.
+4.3 Effects of Demonstration Density
+The relatively low success rates observed in Table 2 are primarily due to our use of keypoint frames in RLBench tasks.
+This approach, widely adopted in the RLBench benchmark [goyal2023rvt, jia2024lift3dfoundationpolicylifting, shridhar2023peract], accelerates both training and evaluation by reducing redundant frames.
+However, selecting only keyframes inevitably decreases the density of demonstration data.
+Since the diffusion policy employed in our imitation learning framework benefits from dense and temporally continuous data due to its long observation horizon, sparse sampling may limit its performance.
+To further investigate demonstration density impact and validate H2R’s robustness, we retrain the diffusion policy with a dense RLBench dataset under identical settings.
+Table 5 shows higher data density significantly boosts overall success rate (e.g., MAE average from 3.3% to 45.0%).
+Critically, H2R consistently enhances performance across sparse and dense data regimes, confirming its general effectiveness regardless of demonstration density.
+Table 5: Effects of Demonstration Density. Success rates (%) across imitation learning tasks in the RLBench environments with sparse data and dense data.
+Data Model RLBench CloseBox CloseLaptopLid ToiletSeatDown Average Sparse data MAE 0.0 10.0 0.0 3.3 MAE+H2R 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) Dense data MAE 50.0 40.0 45.0 45.0 MAE+H2R 60.0 (+10.0%) 45.0 (+5.0%) 60.0 (+15.0%) 55.0 (+10.0%)
 5 Real World Experiment
 Table 6: Real-world Task Results. We report the success rate (%) over real-world tasks for MAE and R3M.
-Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 21.7 43.3 (+21.7%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%)
-5.1 Experiment setup
+Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 23.3 43.3 (+20.0%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%) Franka-PlaceToy 10 15 (+5%) 10 25 (+15%) Franka-PlaceToy 25 40 (+15%) 20 30 (+10%) Franka-SweepRubbish 15 15 (0%) 20 20 (0%) Franka-SweepRubbish 20 35 (+15%) 20 20 (0%) Franka-WeightSauce 40 45 (+5%) 20 30 (+10%) Franka-WeightSauce 30 35 (+25%) 20 25 (+5%) Average 21.7 25 (+3.3%) 17.7 25 (+7.3%) Average 25 36.7 (+16.7%) 20 25 (+5%)
+Table 7: Real-world Success Rates with H2R.
+We report the task success rates (%) using two robot embodiments (UR5, Franka) for data augmentation during pre-training.
+All models are evaluated on downstream tasks using the Leaphand end-effector, under MAE and R3M encoders with DP and ACT policies.
+Despite embodiment mismatch, H2R provides consistent gains over the baseline without augmentation.
+Policy Task MAE (SSv2) MAE (H2R-UR5-SSv2) MAE (H2R-Franka-SSv2) R3M(SSv2) R3M (H2R-UR5-SSv2) R3M (H2R-Franka-SSv2) DP Leaphand-GraspChicken 40 55 (+15%) 35 (-5%) 10 35 (+25%) 20 (+10%) Leaphand-StandCup 35 60 (+25%) 50 (+15%) 20 50 (+30%) 30 (+10%) Leaphand-OpenBox 45 65 (+20%) 45 40 45 (+5%) 45 (+5%) Average 40 60 (+20%) 43.3 (+3.3%) 21.7 43.3 (+21.7%) 31.7 (+10%) ACT Leaphand-GraspChicken 45 50 (+5%) 50 (+5%) 10 35 (+25%) 40 (+30%) Leaphand-StandCup 25 50 (+25%) 50 (+25%) 20 60 (+40%) 25 (+5%) Leaphand-OpenBox 30 40 (+10%) 30 15 20 (+5%) 5 (-10%) Average 33.3 46.7 (+13.3%) 43.3 (+10%) 15 38.3 (+23.3%) 23.3 (+8.3%)
 Robot Setups.
-We validate the effectiveness of H2R in real-world manipulation tasks using a UR5 robotic arm with two different end effectors: Gripper RobotiqAdaptiveGrippers and Leaphand shaw2023leaphandlowcostefficient .
-Realsense IntelRealSenseL515 is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model.
-Real-world setups are shown in Figure 6.
+We evaluate the effectiveness of H2R across three real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [RobotiqAdaptiveGrippers], (ii) a UR5 arm equipped with a Leaphand end effector [shaw2023leaphandlowcostefficient], and (iii) a dual-arm Franka Emika system using parallel Grippers.
+Realsense [IntelRealSenseL515] is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model.
+We provide the visualization of the robot setup in real-world scenarios in the supplementary materials.
 Figure 4: Visualization of Real-world Manipulation Tasks.
 The left columns show Gripper tasks and the right columns show Leaphand tasks.
-Each task is illustrated with six frames, demonstrating the progression from the initial state to the completion of the manipulation.
+Each task is illustrated with three frames, demonstrating the progression from the initial state to the completion of the manipulation.
 Real-world Tasks.
-We set up six tasks for gripper manipulation and dexterous manipulation (Figure 5).
+We set up nine tasks for gripper manipulation, dexterous manipulation and dual-arm manipulation (Figure 4).
 For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed.
 For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened.
-Data Collection.
-We collect expert demonstrations through human teleoperation.
-For gripper manipulation, we use keyboard-based teleoperation.
-For dexterous manipulation, we teleoperate Leaphand using a vision-based retargeting system.
-Due to differences in task complexity, robot embodiment, and teleoperation methods, we adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 4.
+For the dual-arm manipulation tasks, we design: (7) Dualarm-PlaceToy, where the left arm takes a toy from the pink box and the right arm moves it to the blue box; (8) Dualarm-SweepRubbish, where the left arm sweeps the trash into the dustpan held by the right arm; and (9) Dualarm-WeightSauce, where the left arm places the cup on the scale and the right arm picks up the bottle for a pouring motion.
+We adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 8.
 Table 8: Task-specific Data Collection and Evaluation Settings.
-Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800
+Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800 Dualarm-PlaceToy 1000 300 1500 Dualarm-SweepRubbish 1000 300 1500 Dualarm-WeightSauce 1000 300 1500
+Table 9: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
+Policy Task H2R H2R w/o Overlay H2R w/o Retarget DP Leaphand-GraspChicken 55 30 (-25%) 30 (-25%) Leaphand-StandCup 60 40 (-20%) 55 (-5%) Leaphand-OpenBox 65 20 (-45%) 45 (-20%) Average 60 30 (-30%) 43.3 (-16.7%) ACT Leaphand-GraspChicken 50 25 (-25%) 45 (-5%) Leaphand-StandCup 50 35 (-15%) 30 (-20%) Leaphand-OpenBox 40 25 (-15%) 15 (-25%) Average 46.7 28.3 (-18.3%) 30 (-16.7%)
 Policy Training.
-For policy training, we select the Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act as policy frameworks.
-We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.1.
-DP and ACT are trained for 300 epochs for each gripper manipulation task and 3000 epochs for each dexterous manipulation task.
+For policy training, we select the Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] as policy frameworks.
+We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.
 Evaluation.
@@
 We report the success rates of these tasks to assess model quality.
-Gripper Setup
-Leaphand Setup
-Figure 6: Real-world Robot Setups. Experimental setups using a UR5 robot with two different end effectors: a parallel-jaw gripper (left) and a dexterous Leaphand hand (right).
 5.1 Real-world Results
-In the real-world tasks, we employ diffusion policy (DP) chi2023diffusionpolicy and ACT zhao2023act to drive robots in performing real-world tasks, where the visual encoders are pre-trained using MAE and R3M.
-Policy training details are presented in Appendix  C.
-To evaluate the effectiveness and generalization of H2R, we conduct real-world experiments under two different robot embodiments: UR5 (same embodiment as downstream execution) / Franka (cross-embodiment generalization).
-To assess the effectiveness of H2R, we use H2R-UR5-SSv2-1M-Mix during pre-training, which matches the embodiment used in policy training and evaluation.
+In the real-world tasks, we employ diffusion policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] to drive robots in performing manipulation tasks, where the visual encoders are pre-trained using MAE and R3M.
+To evaluate the effectiveness and generalization of H2R, we first conduct experiments to validate its performance enhancement when the pre-training and policy training embodiments are the same.
+We then perform cross-embodiment experiments to investigate the transfer effects of H2R when the pre-training and policy training embodiments differ.
+We finally apply H2R to fine-tune a pre-trained vision backbone in a Vision-Language-Action model, showcasing its promising potential in this advanced setting.
+For identical embodiment experiment, we use UR5 or Franka robot for augmentation during pre-training, which matches the embodiment used in policy training and evaluation.
 As shown in Table 6, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies.
-For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M), while for Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M).
-Similar trends are observed under ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), and Gripper tasks improving by 8.3% and 11.7%, respectively.
-To evaluate cross-embodiment generalization, we use H2R-Franka-SSv2-1M dataset.
-Downstream policy training and evaluation are still conducted with UR5 Leaphand.
+For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M).
+For Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M).
+For Dual-arm Franka tasks, the improvements reach 3.3% (MAE) and 7.3% (R3M).
+Similar trends are observed in ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), Gripper tasks improving by 8.3%(MAE) and 11.7%(R3M), and Dual-arm Franka tasks improving by 16.7%(MAE) and 5%(R3M).
+To evaluate cross-embodiment generalization, we use Franka-augmented the SSv2 dataset with H2R pipeline.
+Downstream policy training and evaluation are conducted with UR5 Leaphand.
 As shown in Table 7, H2R still outperforms raw Ego4D despite the embodiment mismatch.
@@
 Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%.
-Although Franka-based pretraining underperforms UR5, the results confirm H2R’s robustness to embodiment variations.
-Additional analysis, including failure case categorization and experiments on generalization under lighting perturbations, is provided in Appendix  E and  G.
 5.2 Performance on Other Datasets
@@
 Task MAE (Ego4D) MAE (H2R-Ego4D) R3M (Ego4D) R3M (H2R-Ego4D) Leaphand-GraspChicken 25 35 (+10%) 15 20 (+5%) Leaphand-StandCup 25 50 (+25%) 20 25 (+5%) Leaphand-OpenBox 35 45 (+10%) 30 40 (+10%) Average 28.3 43.3 (+15%) 21.7 28.3 (+6.7%)
-We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented H2R-UR5-Ego4D-1M-Mix, and apply them to real-world Leaphand manipulation tasks using the ACT zhao2023act policy.
+We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented dataset, and apply them to real-world Leaphand manipulation tasks using the ACT [zhao2023act] policy.
 As shown in Table 10, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M.
 These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
-Table 6: Real-world Success Rates with H2R.
-We report the task success rates (%) using two robot embodiments (UR5, Franka) for data augmentation during pre-training.
-All models are evaluated on downstream tasks using the Leaphand end-effector, under MAE and R3M encoders with DP and ACT policies.
-Despite embodiment mismatch, H2R provides consistent gains over the baseline without augmentation.
-Policy Task MAE (SSv2) MAE (H2R-UR5-SSv2) MAE (H2R-Franka-SSv2) R3M(SSv2) R3M (H2R-UR5-SSv2) R3M (H2R-Franka-SSv2) DP Leaphand-GraspChicken 40 55 (+15%) 35 (-5%) 10 35 (+25%) 20 (+10%) Leaphand-StandCup 35 60 (+25%) 50 (+15%) 20 50 (+30%) 30 (+10%) Leaphand-OpenBox 45 65 (+20%) 45 40 45 (+5%) 45 (+5%) Average 40 60 (+20%) 43.3 (+3.3%) 21.7 43.3 (+21.7%) 31.7 (+10%) ACT Leaphand-GraspChicken 45 50 (+5%) 50 (+5%) 10 35 (+25%) 40 (+30%) Leaphand-StandCup 25 50 (+25%) 50 (+25%) 20 60 (+40%) 25 (+5%) Leaphand-OpenBox 30 40 (+10%) 30 15 20 (+5%) 5 (-10%) Average 33.3 46.7 (+13.3%) 43.3 (+10%) 15 38.3 (+23.3%) 23.3 (+8.3%)
-Table 7: Ablation Study.
-Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
-Task success rates (%) are reported under DP and ACT policies.
-Removing either component causes significant performance degradation, with red values showing the relative drop compared to full H2R.
-Policy Task H2R H2R w/o Overlay H2R w/o Retarget DP Leaphand-GraspChicken 55 30 (-25%) 30 (-25%) Leaphand-StandCup 60 40 (-20%) 55 (-5%) Leaphand-OpenBox 65 20 (-45%) 45 (-20%) Average 60 30 (-30%) 43.3 (-16.7%) ACT Leaphand-GraspChicken 50 25 (-25%) 45 (-5%) Leaphand-StandCup 50 35 (-15%) 30 (-20%) Leaphand-OpenBox 40 25 (-15%) 15 (-25%) Average 46.7 28.3 (-18.3%) 30 (-16.7%)
+5.3 Potential for VLA Model Pretraining
+Figure 5: Dual-UR5 tasks. The left columns shows PlaceBowl task and the right columns shows StackBlocks task.
+A common paradigm in Vision-Language-Action (VLA) models is to initialize the visual backbone using encoders pre-trained on large-scale datasets (e.g., CLIP, VAEs).
+In this experiment, we investigate whether fine-tuning a pre-trained visual backbone on robot-oriented video data can enhance its suitability for robotic policy learning.
+Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [li2025unifiedvideoactionmodel] separately on both the original SSv2 and the H2R-augmented datasets.
+The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5 robot tasks(shown in Figure  5).
+The results shown in Table  11 demonstrate a clear performance trend in the different variants of the UVA model.
+Table 11: Success rates of UVA variants with different visual backbones.
+Model PlaceBowl StackBlocks Average UVA 0.2 0.2 0.2 UVA + SSv2 0.1 0.2 0.15 UVA + SSv2 + H2R 0.4 0.3 0.35
 5.4 Ablation Study
@@
 H2R enables efficient transfer of task knowledge from human demonstrations to robotic systems, reducing the reliance on costly robot-specific data collection.
-Finally, we expressed the broader impact of this work in Appendix H.
-Appendix A Limitations
-H2R is a simple data augmentation technique for robot pre-training from videos. There are several limitations, which open up possibilities for future work:
-  1. 1.
-    H2R augmentation pipeline is sensitive to the accuracy of pose estimation and hand segmentaion, which demands high-quality egocentric video datasets as prerequisites.
-This sensitivity becomes particularly pronounced in challenging scenarios involving low-light conditions or partial hand occlusions, where current estimation methods often fail to maintain required precision.
-  2. 2.
-    While our robotic platform incorporates diverse manipulator configurations spanning various robotic arm, grippers and dexterous hands, the current implementation remains limited to single-arm operation.
-Future work will expand this framework to dual-arm and humanoid robotics, significantly broadening the model’s application scope.
-  3. 3.
-    It would be valuable to extend our work on more tasks besides manipulation, like visual navigation and mobile manipulation.
-Appendix B Details of Simulator Camera Position Alignment
+[…]
+Supplementary Material
+Appendix A Details of Simulator Camera Position Alignment
 We define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robot arm in the simulator.
@@
 ${}^{W}\mathbf{cam}_{sim}$ $=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$
-Figure 7: H2R samples. Visual comparison between original human data (top) and our augmented data (bottom).
-Appendix C Policy Training Details
-Policy Training in Simulation Experiment.
-For Robomimic tasks, we train for 200 steps and report the mean success rate.
-For tasks in RLBench, we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization.
-For the PushT task, we train the Diffusion Policy model for 200 epoches and report the success rate in the simulation environment.
-The training hyperparameters used in this work are identical to those described in the original paper.
-Policy Training in Real-world Experiment.
-We select the Diffusion Policy (DP) chi2023diffusionpolicy and ACT zhao2023act as policy frameworks.
-We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section  4.1.
-DP and ACT are trained for 300 epochs for each gripper manipulation task and 3000 epochs for each dexterous manipulation task.
-Appendix D Evaluation of H2R Effectiveness
-In Figure 7 are six pairs of subfigures, each showing a human action and its corresponding augmented robotic action. The six action pairs in Figure 7 (left to right) are described as follows:
-  • holding a yellow measuring tape at a construction site
-  • holding a bottle
-  • painting on the wall with a brush / Brushing paint on a wall
-  • turning on the kitchen faucet
-  • preparing painting materials on a table
-The similarity scores computed by CLIP are presented in Table 8.
-Across these samples, the augmented images consistently exhibit higher similarity scores with the robot-centric action descriptions compared to the similarity scores between the unaugmented images and the human-centric action descriptions.
-This quantitative analysis demonstrates that the H2R augmentation effectively enhances the visual alignment between human and robotic actions.
-Table 8: CLIP similarity scores. Higher values indicate better alignment between images and action descriptions.
-Img1 Img2 Img3 Img4 Img5 Img6
-ori 30.6 23.7 31.2 31.7 27.5 28.7
-aug 32.6 28.9 32.3 32.0 28.7 31.0
+Appendix B Details of Experimental Setup
+Figure 6: Simulation benchmark.
+We choose 3 tasks from the Robomimic, 3 tasks from the RLBench, and 3 tasks from the CortexBench, covering a range of robotic manipulation skills.
+We also include the PushT task, designed for the Diffusion Policy framework, as an additional benchmark to evaluate performance in a different task setup.
+Figure 7: Real-world Robot Setup. Illustration of different real-world experimental setup.
+Data Collection in Real-world Experiment.
+For the UR5-gripper, demonstration trajectories were collected via keyboard-based teleoperation.
+For the UR5-Leaphand, demonstrations were gathered using vision-guided teleoperation.
+For the Dual-arm Franka and Dual-arm UR5 setups, data collection was performed through homogeneous-arm bilateral teleoperation.
+Additional Evaluation Details.
+Figure  6 provides visualizations of the different simulation benchmarks, and Figure  7 illustrates the real-world experimental setup.
+For the simulation benchmarks, we use the default configurations for both training and evaluation.
+In the real-world tasks, the UR5–Gripper and UR5–Leaphand setups use the camera shown in Figure  7 as the visual input, whereas the dual-arm Franka and dual-arm UR5 configurations use a front–top camera together with two wrist-mounted cameras (three cameras in total) for visual observation.
+To ensure that the initial conditions are evaluated consistently across algorithms, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstration collection.
+During evaluation, the object placement region is kept identical to that used in data collection, ensuring a fair and consistent comparison.
 Stack
@@
 These images provide qualitative insights into the performance and failure modes of the policy in real deployment, highlighting challenges such as object misalignment, perception noise, and grasp precision.
-Appendix E Failure Case Analysis
+Appendix C Failure Case Analysis
 Table 12: Sub-goal.Task-specific sub-goal evaluation.
@@
 MoveCan Square Lift Average PushT MAE 54 25.5 94.5 58 59.2 MAE+CutMix1 72.0 (+18.0%) 30.0 (+4.5%) 95.0 (+0.5%) 65.7 (+7.7%) 37.5 (-21.7%) MAE+CutMix2 58.0 (+4.0%) 36.0 (+10.5%) 90.0 (-4.5%) 61.3 (+3.3%) 40.0 (-19.2%) MAE+CutMix3 78.0 (+24.0%) 32.0 (+9.3%) 92.0 (-2.5%) 67.3 (+2.7%) 42.0 (-17.2%) MAE+H2R 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) R3M 59.5 20.5 85 55 15 R3M+CutMix1 69.5 (+10.0%) 30.0 (+9.5%) 91.0 (+6.0%) 63.5 (+8.5%) 19.0 (+4.0%) R3M+CutMix2 66.0 (+6.5%) 26.0 (+5.5%) 83.0 (-2.0%) 58.3 (+3.3%) 17.0 (+2.0%) R3M+CutMix3 68.0 (+8.5%) 26.0 (+5.5%) 84.0 (-1.0%) 59.3 (+4.3%) 14.0 (-1.0%) R3M+H2R 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%)
-Appendix F Additional Ablation Study in Simulation Experiment
+Table 14: Generalization under Lighting Variations.
+Success rates (%) under real-world lighting disturbances.
+Models trained with additional lighting augmentations (LightAug) show significant improvements in robustness compared to baseline and H2R-only models.
+Tasks MAE MAE+H2R MAE+H2R+LightAug R3M R3M+H2R R3M+H2R+LightAug Leaphand-GraspChicken 10 10 20 0 0 10 Leaphand-StandCup 20 25 40 5 15 20 Leaphand-OpenBox 5 0 25 10 10 10 Average 11.7 11.7 28.3 5 8.3 13.3
+Appendix D Additional Ablation Study in Simulation Experiment
 In addition to pre-training on the H2R data and raw data, we also applied a simple CutMix baseline to demonstrate the effectiveness of using the robotic arm to cover the human hand, which overlays a fixed set of specific images of robotic arms with grippers onto the original images, ensuring that the overlaid images cover the human hands as much as possible, without exceeding the detected bounding box.
@@
 These results demonstrate the effectiveness of using the robotic arm to cover the human hand in video data, as well as the effectiveness of H2R in imitation learning.
-Table 11: Generalization under Lighting Variations.
-Success rates (%) under real-world lighting disturbances.
-Models trained with additional lighting augmentations (LightAug) show significant improvements in robustness compared to baseline and H2R-only models.
-Tasks MAE MAE+H2R MAE+H2R+LightAug R3M R3M+H2R R3M+H2R+LightAug Leaphand-GraspChicken 10 10 20 0 0 10 Leaphand-StandCup 20 25 40 5 15 20 Leaphand-OpenBox 5 0 25 10 10 10 Average 11.7 11.7 28.3 5 8.3 13.3
-Appendix G Generalization on Light Condition.
+Appendix E Generalization on Light Condition.
 Figure 9: Lighting Setup in Real-World Experiments. Light configuration used in real-world evaluations.
@@
 As shown in Table 14, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.
-Appendix H Broader Impact
-In this work, we proposed a a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric video.
-We focused on the theme above and doesn’t have direct social impact.
-We hope that through our data augmentation methods, more egocentric human datasets can be used for higher-level robot manipulation models, thereby promoting progress in the field of embodied intelligence.

### v4

@@
 Abstract
-Large-scale pre-training using videos has proven effective for robot learning.
+Large-scale pre-training using egocentric human videos has proven effective for robot learning.
 However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots.
-To remedy this, we propose H2R, a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric videos.
+To remedy this, we propose H2R, a human-to-robot data augmentation pipeline that converts egocentric human videos into robot-centric visual data.
+H2R estimates human hand pose from videos, retargets the motion to simulated robotic arms, removes human limbs via segmentation and inpainting, and composites rendered robot embodiments into the original frames with camera-aligned geometry.
 This process explicitly bridges the visual gap between human and robot embodiments during pre-training.
-We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data.
+We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2.
 To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions.
-We validate H2R across four simulation benchmarks—Robomimic, RLBench, PushT and CortexBench, and real-world manipulation tasks with UR5 and Dual-Arm Franka/UR5.
-H2R consistently improves downstream success rates, yielding gains of 1.3%–10.2% in simulation and 3.3%–23.3% in real-world tasks across various visual encoders and policy learning methods.
-We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with VLA models.
+We evaluate H2R through comprehensive experiments in both simulation and real-world settings.
+In simulation, H2R consistently improves downstream success rates across four benchmark suites—Robomimic, RLBench, PushT, and CortexBench—yielding gains of 1.3%–10.2% across different visual encoders and policy learning methods.
+In real-world experiments, H2R improves performance on UR5 and dual-arm Franka/UR5 manipulation platforms, achieving 3.3%–23.3% success rate gains across gripper-based, dexterous, and bimanual tasks.
+We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with vision–language–action models.
 These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains.
-The project page is available in https://sites.google.com/view/h2r-robotics.
-1 Introduction
-Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics.
-Inspired by the remarkable accomplishments of large scale pre-training in computer vision [he2022mae, nair2022r3m, zeng2024learningmanipulationpredictinginteraction, karamcheti2023voltron, datacomp] and natural language processing [devlin2018bert, yu2022pointbert, openai2024gpt4technicalreport, anil2023palm2technicalreport, li2024datacomplm], many efforts have been devoted to harnessing large-scale data to construct generalizable representations in the robotics field [embodimentcollaboration2024openxembodimentroboticlearning, brohan2023rt1roboticstransformerrealworld, brohan2023rt2visionlanguageactionmodelstransfer].
-Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive [duan2023ar2, jang2022bczzeroshottaskgeneralization, brohan2023rt1roboticstransformerrealworld, khazatsky2024droidlargescaleinthewildrobot, fang2023rh20tcomprehensiveroboticdataset, gao2024efficientdatacollectionrobotic, lee2021pickandplacetacklingroboticstacking, herzog2023deeprlscalesorting, kalashnikov2021mtoptcontinuousmultitaskrobotic, duan2024manipulate]; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
-Recent works [nair2022r3m, xiao2022mvp, zeng2024learning] analyze such egocentric human video datasets such as Ego4D [grauman2022ego4d], SSv2 [goyal2017something], and Epic Kitchens [damen2018scaling] with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations.
-However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems.
-Specifically, when utilizing the robot expert data to fine-tune the pre-trained robotic representations for downstream robotic tasks, the model has to learn to bridge the visual gap between the first-person human hand and the robots in addition to acquiring nuanced task-specific skills demonstrated in the robot expert data.
-This would result in increased complexity during the fine-tuning process and suboptimal performance.
+I Introduction
+Pre-training generalizable visual representations is a central challenge in robotic manipulation.
+Recent advances in large-scale pre-training in computer vision and language have significantly improved representation learning across domains [22, 37, 27, 17, 10, 1].
+However, collecting large-scale robot demonstrations remains labor-intensive and costly [13, 29, 14], motivating the use of readily available egocentric human videos as an alternative source for robot pre-training.
+Large-scale egocentric datasets such as Ego4D [20], Something-Something V2 [19], and EPIC-Kitchens [8] capture diverse human-object interactions and have been shown to support transferable visual representations for robotic manipulation.
+However, these datasets are inherently human-centric.
+The visual mismatch between human hands in egocentric videos and robotic embodiments at deployment time introduces a gap that is not explicitly addressed during pre-training, limiting representation transfer.
 To mitigate this issue, we propose H2R (as shown in Figure LABEL:fig:pipeline), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation.
 H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video.
-Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [pavlakos2023reconstructinghands3dtransformers] to accurately detect the position and posture of the human hand in egocentric videos.
+Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [40] to accurately detect the position and posture of the human hand in egocentric videos.
 Then, we simulate the same robot state in simulators to obtain the mask of robot actions.
-In the second stage, we use the Segment Anything Model [kirillov2023segment] to automatically separate human hand from background, and use the inpainting model LaMa [suvorov2021resolutionrobustlargemaskinpainting] to fill the removed hand mask.
+In the second stage, we use the Segment Anything Model [30] to automatically separate human hand from background, and use the inpainting model LaMa [51] to fill the removed hand mask.
 After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video.
@@
 It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
-To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics.
+To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based [41] semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics.
 This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
-To further verify the utility of the augmented datasets, we conduct downstream experiments comparing models pre-trained on original egocentric datasets with those trained on our released H2R-enhanced datasets.
-These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice.
-For the pre-training stage, we apply MAE [he2022mae] and R3M [nair2022r3m] frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video.
-The pretrained visual representations are used in downstream imitation learning pipelines.
-We conduct comprehensive evaluations across both simulation and real-world manipulation tasks.
-In simulation, we evaluate pre-trained visual encoders on four benchmark suites: Robomimic, RLBench, PushT and CortexBench.
-H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, +5.3% for MAE and +7.0% for R3M on the PushT task, and +5.3% for MAE and +1.3% for R3M on the CortexBench task.
-In real-world experiments, we deploy H2R-enhanced encoders on different robot and end-effector settings, evaluated over nine manipulation tasks under two policy frameworks: Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act].
-H2R brings consistent improvements, with average success rate gains of +10.0% (MAE) and +14.1% (R3M) under DP, and +12.8% (MAE) and +13.3% (R3M) under ACT.
-H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively.
-Beyond these evaluations, we also test H2R’s ability to support cross-embodiment transfer by pre-training encoders on Franka data and deploying them on UR robots, indicating that H2R improves representation transfer across different morphologies.
-In addition, fine-tuning VLA backbones on H2R-generated datasets yields markedly higher success rates on dual-arm UR tasks than fine-tuning on human-hand videos or using no fine-tuning.
-These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulation and real-world manipulation settings.
-2 Related Work
+We further evaluate the effectiveness of H2R through a comprehensive set of experiments in both simulation and real-world settings.
+In simulation, we conduct imitation learning experiments on four benchmark suites—Robomimic [36], RLBench [24], PushT [7], and CortexBench [35]— using visual encoders pre-trained with MAE and R3M, and downstream policies trained under standard behavior cloning and diffusion-based policy learning frameworks.
+Across these benchmarks, pre-training with H2R consistently improves downstream success rates, with average gains ranging from 1.3% to 10.2% depending on the encoder and task suite.
+In real-world experiments, we deploy H2R-enhanced visual encoders on multiple robotic platforms, including a UR5 arm with parallel grippers, a UR5 arm with a dexterous Leaphand, and dual-arm systems based on Franka and UR5e.
+Policies are trained using both Diffusion Policy [7] and ACT [57].
+Across gripper-based, dexterous, and bimanual manipulation tasks, H2R yields consistent performance improvements in real-world success rates, with gains ranging from 3.3% to 23.3% across different encoder–policy combinations.
+Beyond these primary evaluations, we conduct additional studies to characterize the generalization properties of H2R.
+These include experiments on pre-training with different egocentric datasets, cross-embodiment transfer where the augmentation embodiment differs from the downstream robot, robustness analyses under varying demonstration densities and lighting conditions, and integration with vision–language–action models through targeted fine-tuning.
+Collectively, these results demonstrate that explicitly reducing the human-to-robot visual gap during pre-training leads to robust and transferable improvements in downstream robotic manipulation performance.
+Figure 1: H2R Pipeline.
+The H2R pipeline includes three steps: (1) hand pose estimation, (2) retargeting to simulate robot arm movements, and (3) removal and inpainting of the human hand to create an augmented image with the robot arm, ensuring seamless integration.
+II Related Work
 Robot Imitation Learning.
-Data-driven policy learning [sergey2018learning, luo2024hilserl, xiao2022mvp, yang2024equibotsim3equivariantdiffusionpolicy, chi2023diffusionpolicy, embodimentcollaboration2024openxembodimentroboticlearning] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation.
-Imitation learning [chi2023diffusionpolicy, ze2024dp3, zhao2023act, mandlekar2021matters] trains policies from successful demonstrations, often supervised by behavior cloning [torabi2018behavioral, florence2021implicit] objectives.
-ACT [zhao2023act] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [chi2023diffusionpolicy, ze2024dp3] are introduced to handle the inherent multimodality of robot motions.
-In addition, CordViP [fu2025cordvip] leverages 3D object-robot correspondences to enhance dexterous manipulation.
+Data-driven policy learning [31, 34, 53, 55, 7, 38] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation.
+Imitation learning [7, 56, 57, 36] trains policies from successful demonstrations, often supervised by behavior cloning [52, 15] objectives.
+ACT [57] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [7, 56] are introduced to handle the inherent multimodality of robot motions.
 Visual Encoder Pretraining for Robotics.
 Visual pretraining improves generalization of robotic policies across diverse tasks.
-Researchers have explored architectural designs [He_2016_CVPR, dosovitskiy12021an], training objectives [He_2022_CVPR, chen2020simple, chen2020big], and dataset compositions [deng2009imagenet, lin2014microsoft, shao2019object365, Schuhmann2021LAION400MOD].
-PVR-Control [parisi22a] shows that pretrained visual representations can outperform direct state-based policies.
-RPT [radosavovic2023robot] tokenizes observations to enable masked prediction pretraining.
-Methods like MVP [radosavovic2022realworld] and R3M [nair2022r3m] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning.
-Voltron [karamcheti2023voltron] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
+Researchers have explored architectural designs [23, 12], training objectives [21, 5, 6], and dataset compositions [9, 33, 48, 47].
+PVR-Control [39] shows that pretrained visual representations can outperform direct state-based policies.
+RPT [44] tokenizes observations to enable masked prediction pretraining.
+Methods like MVP [45] and R3M [37] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning.
+Voltron [27] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
 Cross-Domain Visual Alignment.
 Bridging the domain gap between human and robot visual inputs remains a major challenge.
-WHIRL [bahl2022humantorobotimitationwild] matches task structure from third-person views, while RoVi-Aug [chen2024roviaugrobotviewpointaugmentation] and Mirage [chen2024mirage] manipulate appearance via segmentation or image-space preprocessing.
-EgoMimic [kareer2024egomimicscalingimitationlearning] removes hands and normalizes views to align egocentric perspectives.
-3 H2R: Human-to-Robot Data Augmentation
-In this section, we describe H2R, a data augmentation pipeline for robot learning from egocentric human videos (Figure 1).
-It replaces human hands in every frame with robotic arms equipped with various end effectors, generating a new, visually different dataset.
-This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains.
-We first examine the position and pose of the hand and arm in the image, then map this pose to the robotic arm and end-effector in the simulator.
-Next, we erase the original hand and arm from the image and overlay a rendered image of the simulated robot performing the same pose onto the original image in place of the removed hand and arm.
-Figure 3 illustrates examples of human hand videos that have been processed using H2R.
-3.1 H2R Data Augmentation Pipeline
-Figure 1: H2R Pipeline.
-H2R involves replacing human hands with robotic arms by first using the HaMeR model to detect hand poses and camera parameters.
-The human hand is then removed using the SAM, and the inpainting model LaMa fills in the gap.
-A robot hand is constructed based on the detected pose and keypoints, with the camera perspective adjusted to match the original image.
-Finally, the robot hand is overlaid onto the image, ensuring accurate alignment with the human hand.
-3D Hand Pose Estimation.
-In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information.
-We adopt HaMeR [pavlakos2023reconstructinghands3dtransformers], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal.
-Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
-Robotic Arm and End Effector Construction.
-This step involves building the robotic arm and its end-effectors, which are typically either grippers or dexterous hands.
+WHIRL [2] matches task structure from third-person views, while RoVi-Aug [4] and Mirage [43] manipulate appearance via segmentation or image-space preprocessing.
+EgoMimic [28] removes hands and normalizes views to align egocentric perspectives.
+III H2R: Human-to-Robot Data Augmentation
+In this section, we present H2R, a human-to-robot data augmentation framework that converts egocentric human-hand videos into robot-centric visual observations.
+We detail the full augmentation pipeline, including (1)hand pose estimation, (2)retargeting, and (3)removal and inpainting, and further validate the quality of the generated data using a CLIP-based semantic consistency evaluation.
+An overview of the pipeline is shown in Figure 1.
+III-A H2R Data Augmentation Pipeline
+Step1: Hand Pose Estimation.
+To overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information.
+We adopt HaMeR [40], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image.
+This step outputs the position of the human hand, keypoints, and the intrinsic and extrinsic parameters of the camera.
+The hand position will be used in Step 3 (Removal and Inpainting) to segment the human hand, while the keypoints and camera parameters are used in Step 2 (Retargeting) to adjust the robot’s pose and camera position.
+Step2: Retargeting.
+In this step, we simulate the robot arm to mirror the human hand’s movements detected in Step 1.
+Using the hand keypoints and camera parameters from Step 1, we calculate the joint angles of the robot arm.
+We first build the robotic arm and its end-effectors, which are typically either grippers or dexterous hands.
 For dexterous hands, we estimate joint angles from hand keypoints predicted by HaMeR.
@@
 Since hand keypoints don’t capture the full arm pose (especially for joints unrelated to the hand), we manually set reasonable values for the remaining arm joints to complete a plausible robot configuration.
-Simulator Camera Position Alignment.
-The visual bias introduced by the camera perspective is more significant than the action retargeting itself.
-To address this, we use the hand keypoints and camera parameters from HaMeR to adjust the camera pose in the simulator.
+Next, we use the hand keypoints and camera parameters from HaMeR to adjust the camera pose in the simulator.
 Specifically, we define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robotic arm in the simulator.
@@
 The original camera position $^{W}\mathbf{cam}_{real}$ in the world frame is transformed to the aligned simulator position $^{W}\mathbf{cam}_{sim}$ using transformations from human hand ($^{W}_{H}\mathbf{R}$) and robot simulator ($^{W}_{S}\mathbf{R}$) coordinate systems:
-$^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (1)
-A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix.
-Robot Hand Rendering and Copy-paste.
-After setting the camera, the segmentation mask of the robotic arm is obtained by rendering with the simulator camera.
-We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image.
+$^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\;^{W}_{H}\mathbf{R}^{-1}\;{}^{W}\mathbf{cam}_{real}$ (1)
+Step3: Removal and Inpainting.
+After the robot pose is retargeted in Step 2, we proceed to remove the human hand from the image.
+We use the hand position and keypoints from Step 1 to segment out the human hand and arm regions using the Segment Anything Model (SAM) [30].
+Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [51], a powerful inpainting model, to fill in the removed hand-arm region.
+This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
+Once both the arm segmentation from Step 2 and the inpainted image from Step 3 are available, we overlay the robot arm onto the inpainted image.
+We directly obtain the pixel coordinates of the human hand keypoints from Step 1.
 In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices.
 By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
-Human Arm and Hand Removal.
-After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background.
-We use Segment Anything Model (SAM) [kirillov2023segment] to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR.
-Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [suvorov2021resolutionrobustlargemaskinpainting], a powerful inpainting model, to fill in the removed hand-arm region.
-This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
-3.2 Data Quality Evaluation
-Figure 2: H2R samples. Visual comparison between original human data (top) and our augmented data (bottom).
-To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP [radford2021learningtransferablevisualmodels].
-This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks.
-For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective.
-The human-centric template is defined as: A human is [action]; the robot-centric template is defined as: A robotic arm is [action].
-Here, [action] is a natural language phrase describing the high-level behavior, such as ”holding a bottle” or ”brushing paint on a wall”.
-We use the CLIP ViT-B/32 model to compute the cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt.
-This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description.
-Similarly, we compute the similarity between the original human frame and the human-centric prompt.
-This setup allows us to directly compare semantic alignment before and after augmentation under their respective modalities.
-For each frame, only the action-relevant region is retained, and background context is preserved via inpainting using LaMa.
-We average similarity scores over multiple samples per action class to mitigate variance and isolate the impact of robotic replacement.
-All textual prompts are manually curated to maintain consistency across frames and embodiments.
-Figure 2 contains six pairs of subfigures, each showing a human action and its corresponding augmented robotic action.
-The six action pairs in Figure 2 (left to right) are described as follows: (1) Holding a yellow measuring tape at a construction site; (2) Holding a bottle; (3) Painting on the wall with a brush; (4) Rushing paint on a wall; (5) Turning on the kitchen faucet; (6) Preparing painting materials on a table.
-The CLIP similarity scores are reported in Table 1.
-Across all samples, H2R-augmented images consistently achieve higher similarity with robot-centric action descriptions than original images do with human-centric ones.
-This confirms that H2R augmentation effectively improves visual-semantic alignment between human demonstrations and robotic execution.
-Table 1: CLIP similarity scores. Higher values indicate better alignment between images and action descriptions.
-Img-1 Img-2 Img-3 Img-4 Img-5 Img-6
-Ori 30.6 23.7 31.2 31.7 27.5 28.7
-Aug 32.6 28.9 32.3 32.0 28.7 31.0
-This evaluation provides a lightweight, scalable, and interpretable metric for assessing how well visual hand-to-robot transformations preserve task semantics.
-To quantify data quality, we sample 1,000 image pairs (original and H2R-augmented).
-For each original image, we use Qwen-2.5-VL [bai2025qwen25vltechnicalreport] to generate a verb-noun action description (e.g., “washing bowl”).
-We then form two prompts: “A human is [action]” and “A robot arm is [action]”.
-Using CLIP ViT-B/32, we compute the cosine similarity between: (1) the human prompt and the original image, and (2) the robot prompt and the H2R-augmented image.
-We report the average similarity over all 1,000 samples.
-The raw images achieve an average score of 28.01, while H2R-augmented images reach 29.83.
-This demonstrates that H2R-augmented images exhibit stronger semantic alignment with robot-centric action descriptions, confirming improved visual-semantic coherence for robotic manipulation.
-The results validate our data curation pipeline as effective for generating high-quality, robot-relevant visual data at scale.
-Figure 3: Examples of H2R Augmentation.
-Each column shows images before and after augmentation.
-The top row uses UR5 Leaphand, the middle row uses UR5 Gripper, and the bottom row uses Franka to replace human hands.
-4 Simulation Experiment
-Encoder Pre-training.
-We adopt the MAE [he2022mae, xiao2022mvp] and R3M [nair2022r3m] frameworks for pre-training, each employing a Vision Transformer (ViT) Base [dosovitskiy2021imageworth16x16words] model as the visual encoder.
-The training dataset used is a subset of Ego4D (117K clips) or SSv2 (1M images).
-Both MAE and R3M are pre-trained with $\timesA8$00 GPUs.
-MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate.
-Policy Training.
-We train for 200 steps and report the mean success rate for Robomimic tasks.
-For tasks in RLBench [james2019rlbench], we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization.
-For the PushT task, we train the Diffusion Policy model for 200 epochs and report the success rate in the simulation environment.
-For VC-1 Cortex benchmark, we train policies using behavior cloning composed of a 3-layer MLP for 100 epochs.
-The training hyperparameters used in this work are identical to those described in the original paper.
+III-B Data Quality Evaluation
+We evaluate the visual plausibility and semantic consistency of H2R-augmented data using a vision–language similarity metric based on CLIP [42].
+The evaluation procedure is illustrated in Figure 2, which also shows six representative pairs of original human frames and their corresponding H2R-augmented robot-centric images.
+For each image, we associate a high-level verb–noun action description.
+Specifically, given an original human frame and its H2R-augmented counterpart, we construct two textual prompts describing the same action from different perspectives: a human-centric prompt (“A human is [action]”) and a robot-centric prompt (“A robotic arm is [action]”).
+Image–text cosine similarity is then computed using CLIP ViT-B/32 [42], measuring how well each image aligns with its corresponding semantic description.
+To quantify data quality at scale, we randomly sample 1,000 image pairs from the pre-training dataset.
+For each pair, the action phrase is automatically generated using Qwen-2.5-VL [3].
+We report the average CLIP similarity between original images and human-centric prompts, and between H2R-augmented images and robot-centric prompts.
+The original human images achieve an average similarity score of 28.01, while H2R-augmented images reach 29.83.
+This result indicates that H2R consistently improves semantic alignment with robot-centric action descriptions.
+Overall, the CLIP-based evaluation provides a scalable measure of data quality, supporting the effectiveness of H2R in generating visually and semantically coherent training data for robotic visual pre-training.
+Figure 2: Illustration of the CLIP-based quality evaluation.
+We evaluate augmentation quality by comparing CLIP image–text similarity under human-centric and robot-centric action prompts for original versus H2R-augmented images.
+The results show that H2R improves semantic alignment with robot-centric action descriptions, demonstrating the effectiveness of H2R in generating semantically coherent training data for robotic pre-training.
+IV Experiment
+TABLE I: Summary of Experimental Settings. This table outlines the hyperparameters, datasets, and configurations used for encoder pre-training, simulation benchmarks, and real-world robotic tasks.
+Phase Component Framework Key Configurations
+Encoder Pretraining Method MAE ViT-B, 800 epochs, batch size 128, learning rate $4\times 10^{-4}$, $\timesA8$00 GPUs.
+R3M ViT-B, 20K steps, batch size 256, learning rate $1\times 10^{-4}$, $\timesA8$00 GPUs.
+Dataset SSv2 $\sim$1M frames (subset), original vs. H2R-augmented.
+Ego4D 117K clips subset ($\sim$1M frames), original vs. H2R-augmented.
+Simulation Setup Benchmark Robomimic 200 epochs, 3 tasks.
+RLBench 800 epochs, 3 tasks.
+PushT 200 epochs, 1 task.
+CortexBench 100 epochs, 3 tasks.
+Policy BC 3-layer MLP (Robomimic, CortexBench), default hyperparameters.
+DP Diffusion Policy (PushT, RLBench), default hyperparameters.
+Real-world Setup Scene UR5-Gripper 3 tasks, keyboard-based teleoperation for collection, 1 camera, 30 demos.
+UR5-Leaphand 3 tasks, vision-guided teleoperation for collection, 1 camera, 50 demos.
+Dual-arm Franka 3 tasks, homogeneous-arm bilateral teleoperation for collection, 3 cameras, 300 demos.
+Dual-arm UR5e 2 tasks, homogeneous-arm bilateral teleoperation for collection, 3 cameras, 300 demos.
+Policy DP 300 epochs (UR5-Gripper), 3000 epochs (UR5-Leaphand), 9000 epochs (dual-arm).
+ACT 300 epochs (UR5-Gripper), 3000 epochs (UR5-Leaphand), 9000 epochs (dual-arm).
+UVA Default UVA configuration, 3000 epochs.
+In this section, we conduct a comprehensive experimental evaluation of H2R.
+Experimental configurations are summarized in Table I.
+In simulation, we conduct imitation learning experiments across multiple benchmark suites, including studies on generalization across pre-training datasets, comparisons with robotic-data pre-training, and the effect of demonstration density.
+On real robots, we perform manipulation experiments across multiple platforms and task settings, together with studies on cross-embodiment transfer, compatibility with Vision-Language-Action models, component ablations, and robustness under lighting variations.
+IV-A Experimental Setup
+Visual Representation Pre-training.
+We adopt the MAE [22, 53] and R3M [37] frameworks for visual encoder pre-training, both using a Vision Transformer (ViT-Base) [11] architecture.
+Pre-training is conducted on subsets of SSv2 (1M images) or Ego4D (117K clips).
+For each framework, we consider two settings: pre-training on the original human video data and pre-training with additional H2R augmentation.
+H2R overlays robot embodiments onto human videos using three representative robot types: UR5 with Robotiq Gripper, UR5 with Leaphand, and Franka with Robotiq Gripper.
+Except for cross-embodiment experiments, the robot embodiment used for H2R augmentation is matched to the robot platform of the downstream task.
 Simulation Benchmark.
 For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning.
-Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic [mandlekar2021matters], RLbench [james2019rlbench], and Diffusion Policy [chi2023diffusionpolicy].
+Specifically, we select a total of 10 simulation tasks in different environments, which are from Robomimic [36], RLBench [24], PushT [7] and CortexBench [35].
 In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects.
@@
 We also use three MetaWorld tasks: Assembly, ButtonPress and Hammer from VC-1 Cortex benchmark for evaluation.
-An overview of these simulation tasks is shown in Appendix.
-Table 2: Simulation Benchmark Results.
-Success rates (%) across imitation learning tasks in Robomimic, PushT, RLBench and CortexBench, evaluated with MAE and R3M encoders before and after applying H2R.
-Blue represents an increase in task success rate, while Red represents a decrease.
+Real-world Setup.
+TABLE II: Real-world Manipulation Tasks. Overview of task definitions grouped by robot embodiment.
+Embodiment Task Name Task Description UR5-Gripper PickCube The robot grasps a cube and places it into a bowl.
+Stack The robot stacks a blue cube on top of a yellow cube.
+CloseBox The robot retrieves a cube from a box, places it into a bowl, and closes the box lid.
+UR5-Leaphand GraspChicken The robot grasps a toy chicken and places it into a bowl.
+StandCup The robot grasps a fallen cup and places it upright on the table.
+OpenBox The robot opens an articulated box lid using the dexterous hand.
+Dual-arm Franka PlaceToy The left arm grasps a toy from a pink box, and the right arm transfers it into a blue box.
+SweepRubbish The left arm sweeps trash into a dustpan held by the right arm.
+WeighSauce The left arm places a cup onto a scale, while the right arm performs a pouring motion with a bottle.
+Dual-arm UR5e PlaceBowl The robot places a green bowl on top of a blue bowl.
+StackBlocks The robot stacks blocks in the order of red, yellow, and blue to form a single tower.
+We evaluate the effectiveness of H2R across four real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [46], (ii) a UR5 arm equipped with a Leaphand end effector [49], (iii) a dual-arm Franka Emika system with parallel grippers, and (iv) a dual-arm UR5e platform configured as a human-like bimanual manipulation setup.
+Across these platforms, we evaluate a diverse set of real-world manipulation tasks spanning gripper-based, dexterous, and bimanual manipulation; detailed task definitions are summarized in Table II.
+Demonstration data are collected through human teleoperation, with the collection interface adapted to the manipulation setting.
+Specifically, gripper-based tasks are collected using keyboard-based teleoperation, Leaphand tasks using vision-guided teleoperation following the same setup as in CordViP [16], and dual-arm tasks using homogeneous-arm bilateral teleoperation [54].
+The number of demonstrations varies across setups and task categories, ranging from 30 to 300 episodes depending on task complexity, as summarized in Table I.
+For policy training, we adopt Diffusion Policy (DP) [7] and ACT [57] as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning.
+During evaluation, target objects are randomly initialized within predefined regions following a uniform distribution consistent with expert demonstrations.
+Each task is executed for 20 rollouts, and success rates are reported as the primary evaluation metric.
+IV-B Simulation Results
+Performance on Simulation Benchmarks.
+TABLE III: Simulation Benchmark Results.
+Success rates ($\uparrow$) across diverse imitation learning suites.
+Bold indicates the best performance within each group, and green/red denotes the performance gain/drop after applying H2R.
 All subsequent tables follow the same rule.
-Robomimic Diffusion Policy RLBench CortexBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average Assembly ButtonPress Hammer Average MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) 88.0(+4.0%) 88.0(+8.0%) 100.0 (+4.0%) 92.0(+5.3%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 76.0 56.0 88 73.3 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5.0 (+5.0%) 20.0 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%) 68.0 (-8.0%) 60.0 (+4.0%) 96.0 (+8.0%) 74.7 (+1.3%)
-4.1 Simulation Results
-Table 2 shows encoders trained on augmented data H2R-SSv2 significantly outperform those on original SSv2 across all simulation tasks.
-For Robomimic tasks, MAE achieves a 10.2% average success rate improvement and R3M 6.3%, with MAE seeing a substantial 25.5% gain on MoveCan.
-On PushT, H2R boosts MAE by 5.3% and R3M by 7.0%.
-In RLBench, MAE improves its performance by 10.0% on average and R3M by 5.0%.
-In CortexBench, H2R improves MAE by 5.3% and R3M by 1.3%.
-These results confirm that the H2R-augmented dataset delivers superior visual representations for imitation learning to those learned from large-scale human video datasets like SSv2, as H2R  bridges the visual gap between human hands and robotic arms to enhance vision encoder pre-training performance.
-Table 3: Results on Ego4D Dataset. Success rates (%) across imitation learning tasks in the PushT and RLBench environments.
-Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.6 (+0.0%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
-In addition to SSv2 dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the Ego4D dataset.
-The experimental results in the simulator are shown in Table 3, which indicates that H2R remains highly effective when applied to the Ego4D dataset.
-4.2 Pretraining Model on Robotic Datasets
-Table 4: Comparison with model pretrained on robotic datasets.
-All visual encoder weights are initialized using the R3M pretraining method, with the “R3M-H2R” variant pretrained on the H2R-augmented SSv2 dataset.
-Robomimic MoveCan Square Lift Average R3M 59.5 20.5 85.0 55.0 R3M-H2R 61.5 37.5 85.0 61.3 R3M-DROID 54.0 22.0 96.0 56.7
-To further evaluate how H2R compares to models pretrained directly on robotic datasets, we refer to a recent study [jiang2024robotspretrainrobotsmanipulationcentric], which conducted a systematic comparison between visual representations pretrained on large-scale egocentric human video datasets and those trained on robotic datasets such as DROID [khazatsky2024droidlargescaleinthewildrobot].
-The experimental results, summarized in Table 4, show that models pretrained on robotic datasets often provide limited or even degraded performance on downstream manipulation Robomimic benchmarks.
-Moreover, despite being smaller and simulated rather than real-robot data, the H2R dataset outperforms the DROID dataset in pre-training effectiveness, indicating its strong potential for robotic model pre-training and reflecting its high data quality.
-4.3 Effects of Demonstration Density
-The relatively low success rates observed in Table 2 are primarily due to our use of keypoint frames in RLBench tasks.
-This approach, widely adopted in the RLBench benchmark [goyal2023rvt, jia2024lift3dfoundationpolicylifting, shridhar2023peract], accelerates both training and evaluation by reducing redundant frames.
-However, selecting only keyframes inevitably decreases the density of demonstration data.
-Since the diffusion policy employed in our imitation learning framework benefits from dense and temporally continuous data due to its long observation horizon, sparse sampling may limit its performance.
-To further investigate demonstration density impact and validate H2R’s robustness, we retrain the diffusion policy with a dense RLBench dataset under identical settings.
-Table 5 shows higher data density significantly boosts overall success rate (e.g., MAE average from 3.3% to 45.0%).
-Critically, H2R consistently enhances performance across sparse and dense data regimes, confirming its general effectiveness regardless of demonstration density.
-Table 5: Effects of Demonstration Density. Success rates (%) across imitation learning tasks in the RLBench environments with sparse data and dense data.
-Data Model RLBench CloseBox CloseLaptopLid ToiletSeatDown Average Sparse data MAE 0.0 10.0 0.0 3.3 MAE+H2R 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) Dense data MAE 50.0 40.0 45.0 45.0 MAE+H2R 60.0 (+10.0%) 45.0 (+5.0%) 60.0 (+15.0%) 55.0 (+10.0%)
-5 Real World Experiment
-Table 6: Real-world Task Results. We report the success rate (%) over real-world tasks for MAE and R3M.
-Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 23.3 43.3 (+20.0%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%) Franka-PlaceToy 10 15 (+5%) 10 25 (+15%) Franka-PlaceToy 25 40 (+15%) 20 30 (+10%) Franka-SweepRubbish 15 15 (0%) 20 20 (0%) Franka-SweepRubbish 20 35 (+15%) 20 20 (0%) Franka-WeightSauce 40 45 (+5%) 20 30 (+10%) Franka-WeightSauce 30 35 (+25%) 20 25 (+5%) Average 21.7 25 (+3.3%) 17.7 25 (+7.3%) Average 25 36.7 (+16.7%) 20 25 (+5%)
-Table 7: Real-world Success Rates with H2R.
-We report the task success rates (%) using two robot embodiments (UR5, Franka) for data augmentation during pre-training.
-All models are evaluated on downstream tasks using the Leaphand end-effector, under MAE and R3M encoders with DP and ACT policies.
-Despite embodiment mismatch, H2R provides consistent gains over the baseline without augmentation.
-Policy Task MAE (SSv2) MAE (H2R-UR5-SSv2) MAE (H2R-Franka-SSv2) R3M(SSv2) R3M (H2R-UR5-SSv2) R3M (H2R-Franka-SSv2) DP Leaphand-GraspChicken 40 55 (+15%) 35 (-5%) 10 35 (+25%) 20 (+10%) Leaphand-StandCup 35 60 (+25%) 50 (+15%) 20 50 (+30%) 30 (+10%) Leaphand-OpenBox 45 65 (+20%) 45 40 45 (+5%) 45 (+5%) Average 40 60 (+20%) 43.3 (+3.3%) 21.7 43.3 (+21.7%) 31.7 (+10%) ACT Leaphand-GraspChicken 45 50 (+5%) 50 (+5%) 10 35 (+25%) 40 (+30%) Leaphand-StandCup 25 50 (+25%) 50 (+25%) 20 60 (+40%) 25 (+5%) Leaphand-OpenBox 30 40 (+10%) 30 15 20 (+5%) 5 (-10%) Average 33.3 46.7 (+13.3%) 43.3 (+10%) 15 38.3 (+23.3%) 23.3 (+8.3%)
-Robot Setups.
-We evaluate the effectiveness of H2R across three real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [RobotiqAdaptiveGrippers], (ii) a UR5 arm equipped with a Leaphand end effector [shaw2023leaphandlowcostefficient], and (iii) a dual-arm Franka Emika system using parallel Grippers.
-Realsense [IntelRealSenseL515] is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model.
-We provide the visualization of the robot setup in real-world scenarios in the supplementary materials.
-Figure 4: Visualization of Real-world Manipulation Tasks.
-The left columns show Gripper tasks and the right columns show Leaphand tasks.
-Each task is illustrated with three frames, demonstrating the progression from the initial state to the completion of the manipulation.
-Real-world Tasks.
-We set up nine tasks for gripper manipulation, dexterous manipulation and dual-arm manipulation (Figure 4).
-For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed.
-For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened.
-For the dual-arm manipulation tasks, we design: (7) Dualarm-PlaceToy, where the left arm takes a toy from the pink box and the right arm moves it to the blue box; (8) Dualarm-SweepRubbish, where the left arm sweeps the trash into the dustpan held by the right arm; and (9) Dualarm-WeightSauce, where the left arm places the cup on the scale and the right arm picks up the bottle for a pouring motion.
-We adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 8.
-Table 8: Task-specific Data Collection and Evaluation Settings.
-Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800 Dualarm-PlaceToy 1000 300 1500 Dualarm-SweepRubbish 1000 300 1500 Dualarm-WeightSauce 1000 300 1500
-Table 9: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
-Policy Task H2R H2R w/o Overlay H2R w/o Retarget DP Leaphand-GraspChicken 55 30 (-25%) 30 (-25%) Leaphand-StandCup 60 40 (-20%) 55 (-5%) Leaphand-OpenBox 65 20 (-45%) 45 (-20%) Average 60 30 (-30%) 43.3 (-16.7%) ACT Leaphand-GraspChicken 50 25 (-25%) 45 (-5%) Leaphand-StandCup 50 35 (-15%) 30 (-20%) Leaphand-OpenBox 40 25 (-15%) 15 (-25%) Average 46.7 28.3 (-18.3%) 30 (-16.7%)
-Policy Training.
-For policy training, we select the Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] as policy frameworks.
-We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.
-Evaluation.
-To ensure that the initial conditions are consistent throughout the evaluation, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstrations.
-During the evaluation, each real-world task is rolled out 20 times.
-We report the success rates of these tasks to assess model quality.
-5.1 Real-world Results
-In the real-world tasks, we employ diffusion policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] to drive robots in performing manipulation tasks, where the visual encoders are pre-trained using MAE and R3M.
-To evaluate the effectiveness and generalization of H2R, we first conduct experiments to validate its performance enhancement when the pre-training and policy training embodiments are the same.
-We then perform cross-embodiment experiments to investigate the transfer effects of H2R when the pre-training and policy training embodiments differ.
-We finally apply H2R to fine-tune a pre-trained vision backbone in a Vision-Language-Action model, showcasing its promising potential in this advanced setting.
-For identical embodiment experiment, we use UR5 or Franka robot for augmentation during pre-training, which matches the embodiment used in policy training and evaluation.
-As shown in Table 6, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies.
-For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M).
-For Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M).
-For Dual-arm Franka tasks, the improvements reach 3.3% (MAE) and 7.3% (R3M).
-Similar trends are observed in ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), Gripper tasks improving by 8.3%(MAE) and 11.7%(R3M), and Dual-arm Franka tasks improving by 16.7%(MAE) and 5%(R3M).
-To evaluate cross-embodiment generalization, we use Franka-augmented the SSv2 dataset with H2R pipeline.
-Downstream policy training and evaluation are conducted with UR5 Leaphand.
-As shown in Table 7, H2R still outperforms raw Ego4D despite the embodiment mismatch.
-Under DP, MAE improves from 40% to 43.3%, and R3M from 21.7% to 31.7%.
-Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%.
-5.2 Performance on Other Datasets
-Table 10: Real-world Success Rates with Ego4D Pre-training.
-We report success rates (%) on real-world Leaphand tasks using visual encoders pre-trained on the Ego4D dataset.
-All downstream policies are trained using ACT.
-Task MAE (Ego4D) MAE (H2R-Ego4D) R3M (Ego4D) R3M (H2R-Ego4D) Leaphand-GraspChicken 25 35 (+10%) 15 20 (+5%) Leaphand-StandCup 25 50 (+25%) 20 25 (+5%) Leaphand-OpenBox 35 45 (+10%) 30 40 (+10%) Average 28.3 43.3 (+15%) 21.7 28.3 (+6.7%)
-We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented dataset, and apply them to real-world Leaphand manipulation tasks using the ACT [zhao2023act] policy.
-As shown in Table 10, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M.
-These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
-5.3 Potential for VLA Model Pretraining
-Figure 5: Dual-UR5 tasks. The left columns shows PlaceBowl task and the right columns shows StackBlocks task.
+Method Robomimic PushT RLBench CortexBench MoveCan Square Lift Avg.
+PushT CloseBox CloseLaptopLid ToiletSeatDown Avg.
+Assembly ButtonPress Hammer Avg.
+MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R) 79.5 29.5 95.5 68.2 64.5 5.0 15.0 20.0 13.3 88.0 88.0 100.0 92.0 Gain ($\Delta$) +25.5 +4.0 +1.0 +10.2 +5.3 +5.0 +5.0 +20.0 +10.0 +4.0 +8.0 +4.0 +5.3 R3M (SSv2) 59.5 20.5 85.0 55.0 15.0 0.0 20.0 10.0 10.0 76.0 56.0 88.0 73.3 R3M (H2R) 61.5 37.5 85.0 61.3 22.0 5.0 20.0 20.0 15.0 68.0 60.0 96.0 74.7 Gain ($\Delta$) +2.0 +17.0 0.0 +6.3 +7.0 +5.0 0.0 +10.0 +5.0 -8.0 +4.0 +8.0 +1.3
+Table III shows that visual encoders pre-trained with H2R consistently outperform those trained on the original SSv2 dataset across all evaluated simulation benchmarks for both MAE and R3M visual encoders.
+For Robomimic tasks, H2R leads to clear improvements for both encoders.
+MAE achieves a 10.2% increase in average success rate, while R3M improves by 6.3%.
+Notably, MAE observes a substantial 25.5% gain on the MoveCan task.
+Similar trends are observed on PushT, RLBench, and CortexBench, with consistent performance gains across the evaluated tasks for both MAE and R3M.
+These results demonstrate that H2R augmentation consistently enhances the effectiveness of visual representations learned from large-scale human video datasets for downstream imitation learning across diverse simulation environments.
+Generalization Across Pre-training Datasets.
+TABLE IV: Simulation Results with Ego4D Pre-training. Success rates ($\uparrow$) on PushT and RLBench when pre-training MAE/R3M on an Ego4D subset, comparing original vs. H2R-augmented pre-training.
+Method PushT RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Avg.
+MAE (Ego4D) 51.3 0.0 0.0 5.0 1.7 MAE (H2R) 53.5 10.0 5.0 0.0 5.0 Gain ($\Delta$) +2.2 +10.0 +5.0 -5.0 +3.3 R3M (Ego4D) 13.6 10.0 5.0 5.0 6.7 R3M (H2R) 13.6 15.0 5.0 15.0 11.7 Gain ($\Delta$) 0.0 +5.0 0.0 +10.0 +5.0
+We evaluate H2R using Ego4D to assess its generalization beyond SSv2.
+Specifically, we pre-train both MAE and R3M on an Ego4D subset of comparable scale to SSv2, following the same training protocol as in the previous experiments.
+Table IV reports the resulting imitation learning performance on the PushT task and RLBench benchmarks.
+On PushT, H2R leads to improvement for MAE (+2.2%) while preserving the same performance for R3M.
+On RLBench, H2R improves the average success rate for both encoders, with MAE increasing from 1.7% to 5.0% and R3M from 6.7% to 11.7%.
+These results indicate that the benefits of H2R are not specific to SSv2, and can be consistently transferred to visual representations pre-trained on Ego4D, supporting the general applicability of H2R across different sources of large-scale human video data.
+Pretraining on Robotic Datasets.
+We further compare H2R with visual representations pretrained directly on robotic datasets by including models pretrained on the DROID dataset [29], following the setting reported in the recent study [26].
+All models use the R3M framework for pre-training, where R3M is pretrained on SSv2, R3M-DROID is pretrained on the DROID robotic dataset, and R3M-H2R is pretrained on the H2R-augmented SSv2 dataset.
+We evaluate all representations on the same Robomimic benchmarks.
+As shown in Table V, both robotic data pre-training and H2R augmentation lead to performance improvements over SSv2-only pre-training.
+However, R3M-H2R achieves the strongest overall results.
+Despite DROID being collected from real-world robotic executions, H2R pre-training achieves stronger downstream manipulation performance, suggesting that human-to-robot visual augmentation offers an effective alternative to direct robotic data pretraining.
+TABLE V: Comparison to Robotic-data Pre-training.
+Robomimic success rates ($\uparrow$) for R3M pre-training on SSv2, robotic dataset (DROID), and H2R-augmented SSv2, highlighting H2R as an alternative to direct robotic-data pre-training.
+Method MoveCan Square Lift Avg. R3M 59.5 20.5 85.0 55.0 R3M-DROID 54.0 22.0 96.0 56.7 R3M-H2R 61.5 37.5 85.0 61.3
+TABLE VI: Real-world Task Results.
+Success rates ($\uparrow$) on gripper, dexterous (Leaphand), and bimanual (Franka) tasks using DP and ACT policies with MAE/R3M visual encoders.
+Rows denote pre-training variants, and columns denote tasks.
+(a) DP Policy.
+Policy Method Tasks Gripper Leaphand Franka PickCube Stack CloseBox Avg.
+GraspChicken StandCup OpenBox Avg.
+PlaceToy SweepRubbish WeighSauce Avg.
+DP MAE (SSv2) 45.0 50.0 55.0 50.0 40.0 35.0 45.0 40.0 10.0 15.0 40.0 21.7 MAE (H2R) 65.0 55.0 50.0 56.7 55.0 60.0 65.0 60.0 15.0 15.0 45.0 25.0 Gain ($\Delta$) +20.0 +5.0 -5.0 +6.7 +15.0 +25.0 +20.0 +20.0 +5.0 0.0 +5.0 +3.3 R3M (SSv2) 40.0 55.0 45.0 46.7 10.0 20.0 40.0 23.3 10.0 20.0 20.0 17.7 R3M (H2R) 50.0 70.0 65.0 61.7 35.0 50.0 45.0 43.3 25.0 20.0 30.0 25.0 Gain ($\Delta$) +10.0 +15.0 +20.0 +15.0 +25.0 +30.0 +5.0 +20.0 +15.0 0.0 +10.0 +7.3
+(b) ACT Policy.
+Policy Method Tasks Gripper Leaphand Franka PickCube Stack CloseBox Avg.
+GraspChicken StandCup OpenBox Avg.
+PlaceToy SweepRubbish WeighSauce Avg.
+ACT MAE (SSv2) 25.0 20.0 35.0 26.7 45.0 25.0 30.0 33.3 25.0 20.0 30.0 25.0 MAE (H2R) 30.0 35.0 40.0 35.0 50.0 50.0 40.0 46.7 40.0 35.0 35.0 36.7 Gain ($\Delta$) +5.0 +15.0 +5.0 +8.3 +5.0 +25.0 +10.0 +13.3 +15.0 +15.0 +5.0 +11.7 R3M(SSv2) 25.0 20.0 40.0 28.3 10.0 20.0 15.0 15.0 20.0 20.0 20.0 20.0 R3M (H2R) 30.0 40.0 50.0 40.0 35.0 60.0 20.0 38.3 30.0 20.0 25.0 25.0 Gain ($\Delta$) +5.0 +20.0 +10.0 +11.7 +25.0 +40.0 +5.0 +23.3 +10.0 0.0 +5.0 +5.0
+TABLE VII: Effects of Demonstration Density. Success rates ($\uparrow$) under sparse vs. dense RLBench demonstrations, comparing MAE encoders pre-trained with and without H2R.
+Data Model CloseBox CloseLaptopLid ToiletSeatDown Avg.
+Sparse data MAE 0.0 10.0 0.0 3.3 MAE (H2R) 5.0 15.0 20.0 13.3 Gain ($\Delta$) +5.0 +5.0 +20.0 +10.0 Dense data MAE 50.0 40.0 45.0 45.0 MAE (H2R) 60.0 45.0 60.0 55.0 Gain ($\Delta$) +10.0 +5.0 +15.0 +10.0
+Effects of Demonstration Density.
+We observe relatively low success rates on RLBench under sparse, keypoint-based supervision, a setting commonly adopted in prior RLBench works to improve training efficiency [18, 25, 50].
+This observation motivates an additional study to examine whether the effectiveness of H2R depends on demonstration density, by comparing sparse and dense RLBench training under identical settings.
+As shown in Table VII, increasing demonstration density substantially improves overall task performance (e.g., MAE average success rate increases from 3.3% to 45.0%).
+More importantly, H2R consistently yields performance gains in both sparse and dense regimes.
+This result indicates that H2R is not tailored to a specific data density, but instead provides robust visual representations that remain effective across different demonstration settings.
+IV-C Real-world Results
+Performance on Real-world Manipulation Tasks.
+In real-world experiments, we evaluate H2R on three categories of manipulation tasks-gripper, dexterous, and bimanual—using corresponding robotic platforms, including a UR5 equipped with a Robotiq gripper, a UR5 equipped with a Leaphand end effector, and a dual-arm Franka system.
+We adopt Diffusion Policy (DP) [7] and ACT [57] as policy frameworks, with visual encoders pre-trained using MAE and R3M.
+In this setting, the robot embodiment used for H2R augmentation during pre-training is identical to that used in downstream policy training and evaluation.
+As shown in Table VI, H2R consistently improves real-world success rates across all task categories, encoders, and policies.
+Notably, the most pronounced gains are observed on Leaphand tasks, where H2R improves average performance by 20.0% (MAE, R3M) under DP, and by 13.3% (MAE) and 23.3% (R3M) under ACT.
+Gripper-based and dual-arm Franka tasks also benefit from H2R, with consistent improvements across both MAE and R3M.
+These results demonstrate that H2R effectively enhances real-world manipulation performance.
+TABLE VIII: Real-world results with Ego4D pre-training. Success rates ($\uparrow$) on UR5-Leaphand tasks using ACT, comparing Ego4D pre-training with vs. without H2R augmentation.
+Method GraspChicken StandCup OpenBox Avg.
+MAE (Ego4D) 25.0 25.0 35.0 28.3 MAE (H2R) 35.0 50.0 45.0 43.3 Gain ($\Delta$) +10.0 +25.0 +10.0 +15.0 R3M (Ego4D) 15.0 20.0 30.0 21.7 R3M (H2R) 20.0 25.0 40.0 28.3 Gain ($\Delta$) +5.0 +5.0 +10.0 +6.7
+TABLE IX: Cross-Embodiment Real-world Results.
+UR5-Leaphand task success rates ($\uparrow$) when the robot embodiment used for H2R augmentation during pre-training differs from the downstream embodiment (UR5-based vs.
+Franka-based augmentation).
+(a) DP Policy.
+Policy Method GraspChicken StandCup OpenBox Avg.
+DP MAE (SSv2) 40.0 35.0 45.0 40.0 MAE (H2R-UR5) 55.0 60.0 65.0 60.0 Gain ($\Delta$) +15.0 +25.0 +20.0 +20.0 MAE (H2R-Franka) 35.0 50.0 45.0 43.3 Gain ($\Delta$) -5.0 +15.0 0.0 +3.3 R3M (SSv2) 10.0 20.0 40.0 23.3 R3M(H2R-UR5) 35.0 50.0 45.0 43.3 Gain ($\Delta$) +25.0 +30.0 +5.0 +20.0 R3M(H2R-Franka) 20.0 30.0 45.0 31.7 Gain ($\Delta$) +10.0 +10.0 +5.0 +10.0
+(b) ACT Policy.
+Policy Method GraspChicken StandCup OpenBox Avg.
+ACT MAE (SSv2) 45.0 25.0 30.0 33.3 MAE (H2R-UR5) 50.0 50.0 40.0 46.7 Gain ($\Delta$) +5.0 +25.0 +10.0 +13.3 MAE (H2R-Franka) 50.0 50.0 30.0 43.3 Gain ($\Delta$) +5.0 +25.0 0.0 +10.0 R3M (SSv2) 10.0 20.0 15.0 15.0 R3M (H2R-UR5) 35.0 60.0 20.0 38.3 Gain ($\Delta$) +25.0 +40.0 +5.0 +23.3 R3M (H2R-Franka) 40.0 25.0 5.0 23.3 Gain ($\Delta$) +30.0 +5.0 -10.0 +8.3
+Generalization Across Pre-training Datasets.
+To examine whether the effectiveness of H2R observed in simulation also extends to real-world settings, we further evaluate visual encoders pre-trained on Ego4D in real-world Leaphand manipulation tasks.
+Following the same pre-training protocol as in the simulation experiments, both MAE and R3M are trained on an Ego4D subset of comparable scale, with and without H2R augmentation, and deployed using the ACT policy.
+Table VIII reports the success rates across three real-world Leaphand tasks.
+For MAE, applying H2R consistently improves performance on all tasks, increasing the average success rate from 28.3% to 43.3%.
+For R3M, H2R also yields consistent gains, with the average success rate improving from 21.7% to 28.3%.
+These results are consistent with the trends observed in simulation and indicate that the benefits of H2R extend to real-world manipulation when visual encoders are pre-trained on Ego4D.
+Cross-embodiment Generalization.
+We further evaluate the cross-embodiment generalization of H2R by applying H2R-augmented pre-training using a robot embodiment that differs from the one used in downstream policy training and evaluation.
+Specifically, we consider two augmentation settings during pre-training: UR5-based H2R and Franka-based H2R, while all downstream policies are trained and evaluated on UR5-Leaphand tasks.
+As shown in Table IX, H2R continues to improve performance under embodiment mismatch for both MAE and R3M encoders, as well as for both DP and ACT policies.
+Using UR5-based H2R augmentation yields the strongest gains, with average success rates increasing by 20.0% (MAE) and 21.7% (R3M) under DP, and by 13.3% (MAE) and 23.3% (R3M) under ACT.
+When H2R is applied using Franka-based augmentation, performance improvements remain consistent but are generally smaller in magnitude.
+These results suggest that the effectiveness of H2R is not strictly limited to cases where the pre-training and downstream embodiments are identical.
+Within the evaluated settings, H2R remains beneficial even when the augmentation embodiment differs from that used in policy learning, indicating a certain degree of robustness to embodiment variation.
+Synergy with Vision-Language-Action Models.
 A common paradigm in Vision-Language-Action (VLA) models is to initialize the visual backbone using encoders pre-trained on large-scale datasets (e.g., CLIP, VAEs).
 In this experiment, we investigate whether fine-tuning a pre-trained visual backbone on robot-oriented video data can enhance its suitability for robotic policy learning.
-Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [li2025unifiedvideoactionmodel] separately on both the original SSv2 and the H2R-augmented datasets.
-The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5 robot tasks(shown in Figure  5).
-The results shown in Table  11 demonstrate a clear performance trend in the different variants of the UVA model.
-Table 11: Success rates of UVA variants with different visual backbones.
-Model PlaceBowl StackBlocks Average UVA 0.2 0.2 0.2 UVA + SSv2 0.1 0.2 0.15 UVA + SSv2 + H2R 0.4 0.3 0.35
-5.4 Ablation Study
+Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [32] separately on both the original SSv2 and the H2R-augmented datasets.
+The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5e robot tasks.
+The results shown in Table  X demonstrate a clear performance trend in the different variants of the UVA model.
+TABLE X: UVA with different visual backbones. Success rates ($\uparrow$) on dual-arm UR5e tasks when using UVA with original backbone, and with VAEs fine-tuned on SSv2 or H2R-augmented data.
+Method PlaceBowl StackBlocks Avg. UVA 0.20 0.20 0.20 UVA (SSv2) 0.10 0.20 0.15 UVA (H2R) 0.40 0.30 0.35
+TABLE XI: Ablation Study. Ablation results by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
+Policy Method GraspChicken StandCup OpenBox Avg.
+DP H2R 55.0 60.0 65.0 60.0 H2R w/o Overlay 30.0 40.0 20.0 30.0 Gain ($\Delta$) -25.0 -20.0 -45.0 -30.0 H2R w/o Retarget 30.0 55.0 45.0 43.3 Gain ($\Delta$) -25.0 -5.0 -20.0 -16.7 ACT H2R 50.0 50.0 40.0 46.7 H2R w/o Overlay 25.0 35.0 25.0 28.3 Gain ($\Delta$) -25.0 -15.0 -15.0 -18.3 H2R w/o Retarget 45.0 30.0 15.0 30.0 Gain ($\Delta$) -5.0 -20.0 -25.0 -16.7
+Ablation Study.
 To evaluate the effectiveness of each component in H2R, we conduct ablation studies on two time-consuming steps: (1) performing hand inpainting without overlaying a robotic arm (H2R w/o Overlay), and (2) overlaying the arm without precise alignment between the hand and the camera, instead using random pasting (H2R w/o Retarget).
-Table 9 shows the necessity and effectiveness of each component in H2R.
+Table XI shows the necessity and effectiveness of each component in H2R.
 The first step leads to a significant drop in success rate due to the loss of critical human-object interaction pixels after inpainting.
 The second step fails to provide accurate motion cues for the model and introduces visual mismatches with real-world manipulation tasks.
-6 Conclusion
+TABLE XII: Generalization under Lighting Variations.
+Success rates ($\uparrow$) on UR5-Leaphand tasks under real-world lighting disturbances, comparing no augmentation, H2R, and H2R with lighting augmentation (H2R+LightAug).
+Method GraspChicken StandCup OpenBox Avg.
+MAE 10.0 20.0 5.0 11.7 MAE (H2R) 10.0 25.0 0.0 11.7 Gain ($\Delta$) 0.0 +5.0 -5.0 0.0 MAE (H2R + LightAug) 20.0 40.0 25.0 28.3 Gain ($\Delta$) +10.0 +20.0 +20.0 +16.6 R3M 0.0 5.0 10.0 5.0 R3M (H2R) 0.0 15.0 10.0 8.3 Gain ($\Delta$) 0.0 +10.0 0.0 +3.3 R3M (H2R + LightAug) 10.0 20.0 10.0 13.3 Gain ($\Delta$) +10.0 +15.0 0.0 +8.3
+Robustness to Lighting Variations.
+To evaluate the generalization under varying lighting conditions, we introduce illumination disturbances during evaluation.
+Additionally, during training, we incorporate randomized lighting with varying directions and colors into the simulation environment for data augmentation.
+We compare three settings: no augmentation(MAE, R3M), H2R augmentation(H2R), and H2R with lighting disturbances(H2R+LightAug).
+As shown in Table XII, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.
+V Conclusion
 We propose H2R, a data augmentation technique that bridges the visual gap between human hand demonstrations and robotic arm manipulations by replacing human hands in first-person videos with robotic arm movements.
@@
 H2R enables efficient transfer of task knowledge from human demonstrations to robotic systems, reducing the reliance on costly robot-specific data collection.
-[…]
-Supplementary Material
-Appendix A Details of Simulator Camera Position Alignment
-We define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robot arm in the simulator.
-We build the coordinate system $^{W}\mathbf{I}_{H}$ based on the hand keypoints:
-$^{W}\mathbf{I}_{H}=\{^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}\}$ (2)
-Where $^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}$ are unit vectors along the x-axes, y-axes and z-axes of the human hand coorinate system.
-With the keypoints get in HaMeR, we build the three axis of coordinates with the following functions:
-${}^{w}\mathbf{i}_{H,x}$ $=^{w}\mathbf{i}_{0,9}$ (3)
-${}^{w}\mathbf{i}_{H,y}$ $=^{w}\mathbf{i}_{0,9}\times^{w}\mathbf{i}_{0,13}$
-${}^{w}\mathbf{i}_{H,z}$ $=^{w}\mathbf{i}_{H,x}\times^{w}\mathbf{i}_{H,y}$
-Where $^{w}\mathbf{i}_{0,9}$ and $^{w}\mathbf{i}_{0,13}$ are unit vectors along the middle and ring fingers, respectively.
-In this notation, the first index (0) refers to the specific finger (middle or ring), and the second index (9 and 13) corresponds to the joint numbers along those fingers, as defined by the MANO model.
-Similarly, To construct the mapping from hand pose to robot arms, we need to get another coordinate system $^{W}\mathbf{I}_{S}$ in the simulator:
-$^{W}\mathbf{I}_{S}=\{^{w}\mathbf{i}_{S,x},^{w}\mathbf{i}_{S,y},^{w}\mathbf{i}_{S,z}\}$ (4)
-The method of determining the axis of coordinates is the same:
-${}^{w}\mathbf{i}_{S,x}$ $=^{w}\mathbf{i}_{0,2}$ (5)
-${}^{w}\mathbf{i}_{S,y}$ $=^{w}\mathbf{i}_{0,2}\times^{w}\mathbf{i}_{0,3}$
-${}^{w}\mathbf{i}_{s,z}$ $=^{w}\mathbf{i}_{S,x}\times^{w}\mathbf{i}_{S,y}$
-Where $\mathbf{i}_{0,2},\mathbf{i}_{0,3}$ are unit vectors along robot fingers that correspond to human middle and ring fingers and the index corresponds to the joint numbers defined by MANO.
-We build the following two coordinate transformation matrix to construct the mapping:
-${}^{W}_{H}\mathbf{R}$ $=\begin{pmatrix}{}^{W}\mathbf{I}_{H}&\mathbf{key}_{human}\\ \mathbf{O}&1\end{pmatrix}$ (6)
-${}^{W}_{S}\mathbf{R}$ $=\begin{pmatrix}{{}^{W}\mathbf{I}_{S}}&{\mathbf{key}_{robot}}\\ \mathbf{O}&{1}\end{pmatrix}$
-Where $\mathbf{key}_{human},\mathbf{key_{robot}}$ are the positions of human wrist and robot wrist.
-After obtaining the two coordinate systems, we need to determine the position of the camera in the simulator ($^{W}\mathbf{cam}_{sim}$) and the position of the camera in the real world ($^{H}\mathbf{cam}_{Real}$), thus we can ensure we get the same pose of the human hand and robot arms
-${}^{H}\mathbf{cam}_{Real}$ $=^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (7)
-${}^{S}\mathbf{cam}_{sim}$ $=^{H}\mathbf{cam}_{Real}$
-${}^{W}\mathbf{cam}_{sim}$ $=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$
-Appendix B Details of Experimental Setup
-Figure 6: Simulation benchmark.
-We choose 3 tasks from the Robomimic, 3 tasks from the RLBench, and 3 tasks from the CortexBench, covering a range of robotic manipulation skills.
-We also include the PushT task, designed for the Diffusion Policy framework, as an additional benchmark to evaluate performance in a different task setup.
+Appendix A Details of Experimental Setup
 Figure 4: Real-world Robot Setup. Illustration of different real-world experimental setup.
-Data Collection in Real-world Experiment.
-For the UR5-gripper, demonstration trajectories were collected via keyboard-based teleoperation.
-For the UR5-Leaphand, demonstrations were gathered using vision-guided teleoperation.
-For the Dual-arm Franka and Dual-arm UR5 setups, data collection was performed through homogeneous-arm bilateral teleoperation.
-Additional Evaluation Details.
-Figure  6 provides visualizations of the different simulation benchmarks, and Figure  7 illustrates the real-world experimental setup.
-For the simulation benchmarks, we use the default configurations for both training and evaluation.
-In the real-world tasks, the UR5–Gripper and UR5–Leaphand setups use the camera shown in Figure  7 as the visual input, whereas the dual-arm Franka and dual-arm UR5 configurations use a front–top camera together with two wrist-mounted cameras (three cameras in total) for visual observation.
-To ensure that the initial conditions are evaluated consistently across algorithms, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstration collection.
-During evaluation, the object placement region is kept identical to that used in data collection, ensuring a fair and consistent comparison.
+TABLE XIII: Robomimic Experiment result. We report the success rate (%) over IL-based tasks for MAE and R3M Robomimic.
+MoveCan Square Lift Average PushT MAE 54 25.5 94.5 58 59.2 MAE+CutMix1 72.0 (+18.0%) 30.0 (+4.5%) 95.0 (+0.5%) 65.7 (+7.7%) 37.5 (-21.7%) MAE+CutMix2 58.0 (+4.0%) 36.0 (+10.5%) 90.0 (-4.5%) 61.3 (+3.3%) 40.0 (-19.2%) MAE+CutMix3 78.0 (+24.0%) 32.0 (+9.3%) 92.0 (-2.5%) 67.3 (+2.7%) 42.0 (-17.2%) MAE+H2R 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) R3M 59.5 20.5 85 55 15 R3M+CutMix1 69.5 (+10.0%) 30.0 (+9.5%) 91.0 (+6.0%) 63.5 (+8.5%) 19.0 (+4.0%) R3M+CutMix2 66.0 (+6.5%) 26.0 (+5.5%) 83.0 (-2.0%) 58.3 (+3.3%) 17.0 (+2.0%) R3M+CutMix3 68.0 (+8.5%) 26.0 (+5.5%) 84.0 (-1.0%) 59.3 (+4.3%) 14.0 (-1.0%) R3M+H2R 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%)
+A-A Additional Evaluation Details.
+Figure  3 provides visualizations of the different simulation benchmarks, and Figure  4 illustrates the real-world experimental setup.
 Stack
@@
 These images provide qualitative insights into the performance and failure modes of the policy in real deployment, highlighting challenges such as object misalignment, perception noise, and grasp precision.
-Appendix C Failure Case Analysis
-Table 12: Sub-goal.Task-specific sub-goal evaluation.
+Appendix B Failure Case Analysis
+TABLE XIV: Task-specific sub-goal evaluation.
 To gain fine-grained insights into policy performance, we design a manual rubric covering key sub-goals for each manipulation task.
@@
 To enable fine-grained evaluation of policy performance and gain deeper insights into failure cases, we designed a task-specific evaluation rubric.
-Table 12 displays our rubric that the evaluator filled out when rolling out different policies.
-Take the DP policy as an example, the results in Table 12 demonstrate that H2R-augmented visual representation models not only improve overall success rates in real-world tasks, but also allow to accomplish more than half of the task consistently.
-Table 13: Robomimic Experiment result. We report the success rate (%) over IL-based tasks for MAE and R3M Robomimic.
-MoveCan Square Lift Average PushT MAE 54 25.5 94.5 58 59.2 MAE+CutMix1 72.0 (+18.0%) 30.0 (+4.5%) 95.0 (+0.5%) 65.7 (+7.7%) 37.5 (-21.7%) MAE+CutMix2 58.0 (+4.0%) 36.0 (+10.5%) 90.0 (-4.5%) 61.3 (+3.3%) 40.0 (-19.2%) MAE+CutMix3 78.0 (+24.0%) 32.0 (+9.3%) 92.0 (-2.5%) 67.3 (+2.7%) 42.0 (-17.2%) MAE+H2R 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) R3M 59.5 20.5 85 55 15 R3M+CutMix1 69.5 (+10.0%) 30.0 (+9.5%) 91.0 (+6.0%) 63.5 (+8.5%) 19.0 (+4.0%) R3M+CutMix2 66.0 (+6.5%) 26.0 (+5.5%) 83.0 (-2.0%) 58.3 (+3.3%) 17.0 (+2.0%) R3M+CutMix3 68.0 (+8.5%) 26.0 (+5.5%) 84.0 (-1.0%) 59.3 (+4.3%) 14.0 (-1.0%) R3M+H2R 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%)
-Table 14: Generalization under Lighting Variations.
-Success rates (%) under real-world lighting disturbances.
-Models trained with additional lighting augmentations (LightAug) show significant improvements in robustness compared to baseline and H2R-only models.
-Tasks MAE MAE+H2R MAE+H2R+LightAug R3M R3M+H2R R3M+H2R+LightAug Leaphand-GraspChicken 10 10 20 0 0 10 Leaphand-StandCup 20 25 40 5 15 20 Leaphand-OpenBox 5 0 25 10 10 10 Average 11.7 11.7 28.3 5 8.3 13.3
-Appendix D Additional Ablation Study in Simulation Experiment
+Table XIV displays our rubric that the evaluator filled out when rolling out different policies.
+Take the DP policy as an example, the results in Table XIV demonstrate that H2R-augmented visual representation models not only improve overall success rates in real-world tasks, but also allow to accomplish more than half of the task consistently.
+Appendix C Additional Ablation Study in Simulation Experiment
 In addition to pre-training on the H2R data and raw data, we also applied a simple CutMix baseline to demonstrate the effectiveness of using the robotic arm to cover the human hand, which overlays a fixed set of specific images of robotic arms with grippers onto the original images, ensuring that the overlaid images cover the human hands as much as possible, without exceeding the detected bounding box.
@@
 Based on the type of robotic arm used in CutMix, we categorize the augmented set into three types: CutMix1 represents the UR5 robotic arm, CutMix2 refers to the Franka robotic arm, and CutMix3 combines both the UR5 and Franka robotic arms.
-From Table 13, we observe that the encoder trained on H2R processed data shows consistent improvements across various tasks compared to the encoder trained on the original data, with the average success rate on all tasks ranging from 0.9% to 10.2%.
+From Table XIII, we observe that the encoder trained on H2R processed data shows consistent improvements across various tasks compared to the encoder trained on the original data, with the average success rate on all tasks ranging from 0.9% to 10.2%.
 Especially for the more challenging MoveCan task, it can improve the success rate by 25.5%.
@@
 These results demonstrate the effectiveness of using the robotic arm to cover the human hand in video data, as well as the effectiveness of H2R in imitation learning.
-Appendix E Generalization on Light Condition.
-Figure 9: Lighting Setup in Real-World Experiments. Light configuration used in real-world evaluations.
-To evaluate the generalization under varying lighting conditions, we introduce illumination disturbances during evaluation, as illustrated in Figure 9.
-Additionally, during training, we incorporate randomized lighting with varying directions and colors into the simulation environment for data augmentation.
-We compare three settings: no augmentation(MAE, R3M), H2R augmentation(MAE+H2R, R3M+H2R), and H2R with lighting disturbances(MAE+H2R+LightAug, R3M+H2R+LightAug).
-As shown in Table 14, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.